### v1

# Revealing POMDPs: Qualitative and Quantitative Analysis for Parity Objectives

Abstract
Partially observable Markov decision processes (POMDPs) are a central model for uncertainty in sequential decision making. The most basic objective is the reachability objective, where a target set must be eventually visited, and the more general parity objectives can model all $\omega$-regular specifications. For such objectives, the computational analysis problems are the following: (a) qualitative analysis that asks whether the objective can be satisfied with probability $1$ (almost-sure winning) or probability arbitrarily close to $1$ (limit-sure winning); and (b) quantitative analysis that asks for the approximation of the optimal probability of satisfying the objective. For general POMDPs, almost-sure analysis for reachability objectives is EXPTIME-complete, but limit-sure and quantitative analyses for reachability objectives are undecidable; almost-sure, limit-sure, and quantitative analyses for parity objectives are all undecidable. A special class of POMDPs, called revealing POMDPs, has been studied recently in several works, and for this subclass the almost-sure analysis for parity objectives was shown to be EXPTIME-complete. In this work, we show that for revealing POMDPs the limit-sure analysis for parity objectives is EXPTIME-complete, and even the quantitative analysis for parity objectives can be achieved in EXPTIME.
1 Introduction
POMDPs
Partially observable Markov decision processes (POMDPs) model sequential decision-making under uncertainty (Bertsekas 1976; Papadimitriou and Tsitsiklis 1987; Kaelbling, Littman, and Cassandra 1998). At each step, the environment is in some hidden state. A controller interacts with it by choosing actions. The chosen action and the current state determine a probability distribution over the subsequent state. The controller cannot observe the state directly, but observes a signal that discloses only partial information of the state. POMDPs generalize classic models: Markov decision processes (MDPs), in which the state is fully observed (Puterman 2014), and blind MDPs, in which no state information is observed and are equivalent to Probabilistic Finite Automata (Rabin 1963; Paz 1971).
Objectives
The controller aims to maximize an objective function, which formally captures the desired behaviors of the model. Two main classes of objectives are typically considered: logical objectives, e.g., reachability and LTL (linear-temporal logic), and quantitative objectives, e.g., discounted sum and limit-average; see (Puterman 2014; Baier and Katoen 2008; Filar and Vrieze 1997) for details. This work focuses on logical objectives.
The most basic example of logical objectives is reachability, i.e., given a set of target states, the objective requires that some target state is visited at least once. A more general class of logical objectives is parity objectives, which assign to each state a non-negative integer, called priority. The objective is satisfied if the smallest priority appearing infinitely often is even. Parity objectives express all commonly used temporal objectives such as liveness, and are a canonical form to express all $\omega$-regular objectives (Thomas 1997) and LTL objectives (Baier and Katoen 2008). Hence, the study of POMDPs with parity objectives is a fundamental theoretical problem.
Applications
POMDPs have been applied across diverse fields, including computational biology (Durbin et al. 1998) and reinforcement learning (Kaelbling, Littman, and Moore 1996). In particular, POMDPs with logical objectives have proven useful in many application areas such as probabilistic planning (Bonet and Geffner 2009); randomized embedded scheduler (de Alfaro et al. 2005); randomized distributed algorithms (Pogosyants, Segala, and Lynch 2000); probabilistic specification languages (Baier, Größer, and Bertrand 2012); and robot planning (Kress-Gazit, Fainekos, and Pappas 2009; Chatterjee et al. 2015).
Computational Analysis Questions
A policy determines the choice of actions by the controller. The value corresponds to the maximum probability that the controller can guarantee to satisfy an objective. The main computational analysis of POMDPs with reachability and parity objectives considers the following two problems.
  • (a)
    Qualitative analysis has two variants: the almost-sure winning asks whether there exists a policy that satisfies the objective with probability $1$; and the limit-sure winning asks whether the objective can be satisfied with probability arbitrarily close to $1$.
  • (b)
    Quantitative analysis asks to approximate the maximum or optimal probability with which the objective can be satisfied, up to a given additive error.
General Undecidability
Most results for POMDPs and the computational analysis are negative (undecidability results). Indeed, the limit-sure analysis for reachability objectives is undecidable (Gimbert and Oualhadj 2010; Chatterjee and Henzinger 2010), which extends to general parity objectives. The almost-sure analysis for reachability objectives is EXPTIME-complete (Baier, Bertrand, and Größer 2008; Chatterjee, Doyen, and Henzinger 2010), but is undecidable for parity objectives even for two priorities (namely coBüchi objectives) (Baier, Bertrand, and Größer 2008; Chatterjee et al. 2010). The quantitative analysis for reachability objectives is undecidable (Madani, Hanks, and Condon 2003), which extends to general parity objectives. Given this wide range of results about non-existence of algorithms a natural direction to explore is the existence of subclasses for which the computational problems are decidable.
Revealing POMDPs
We study revealing POMDPs, a special subclass of POMDPs where each visited state is announced to the controller with a positive probability. Consequently, the controller’s uncertainty about the state, i.e. the probability distribution over states (the belief), occasionally collapses into a Dirac distribution on the announced state. Revealing POMDPs have been previously studied in several works, including (Chen and Liew 2023; Belly et al. 2025; Avrachenkov, Dhiman, and Kavitha 2025), which present motivation and applications for this model.
Previous Results and Open Questions
A key result for revealing POMDPs shows that the almost-sure analysis for parity objectives is EXPTIME-complete (Belly et al. 2025). However, the limit-sure and quantitative analysis problems for reachability and parity objectives remained open for revealing POMDPs.
Our Contributions
We address the open questions for revealing POMDPs. Our main contributions are the following. For revealing POMDPs with parity objectives, we show that
  • The limit-sure analysis coincides with almost-sure analysis, and consequently is EXPTIME-complete.
  • The quantitative analysis can be achieved in EXPTIME, i.e., in the same complexity as for qualitative analysis.
The results for POMDPs and revealing POMDPs are summarized in Table 1 and Table 2, respectively.
Problems Objectives
Reachability Parity
Almost-sure EXP-complete Undecidable
Limit-sure Undecidable Undecidable
Quantitative Undecidable Undecidable
Table 1: Computational complexity for POMDPs with reachability and parity objectives.
Problems Objectives
Reachability Parity
Almost-sure EXP EXP-complete
Limit-sure EXP EXP-complete
Quantitative EXP EXP
Table 2: Computational complexity for revealing POMDPs with reachability and parity objectives. Our contributions are marked in bold.
Technical Contributions
A closely related work established that almost-sure analysis for parity objectives in revealing POMDPs is EXPTIME-complete (Belly et al. 2025). Additionally, (Chatterjee, Chmelík, and Tracol 2016) previously showed that almost-sure analysis for parity objectives in general POMDPs under finite-memory policies is EXPTIME-complete. Therefore, the EXPTIME-completeness result naturally follows once finite memory policies are proven sufficient for almost-sure analysis. However, limit-sure analysis and quantitative analysis for POMDPs remain undecidable in general, even under finite memory policies (Chatterjee, Saona, and Ziliotto 2021). This highlights a sharp contrast with the almost-sure analysis and gives rise to new technical challenges.
First, we consider revealing POMDPs with the belief-reachability objectives. Given a set of target states, the belief-reachability objectives consider the probability of eventually observing such states. We prove that the quantitative analysis of belief-reachability in revealing POMDPs can be achieved in EXPTIME. The argument proceeds in two steps: $(i)$ we prove that this objective can be approximated by their finite-horizon counterparts; and $(ii)$ we show that this finite-horizon objective can be approximated by a point-based approximation.
Then, we consider revealing POMDPs with the parity objectives. We show that the value for parity objectives corresponds to the value for belief-reachability objectives to almost-sure winning states. Finally, we prove the existence of optimal policies for parity objectives in revealing POMDPs.
Proofs and details omitted due to space restrictions are provided in the Appendix.
Related Works
The study of subclasses of POMDPs with tractable algorithms is a broad topic with many directions. For example, subclasses of POMDPs for qualitative analysis have been explored in (Fijalkow et al. 2015; Chatterjee and Tracol 2012); and for quantitative analysis various subclasses have also been studied such as with ergodicity condition (Chatterjee et al. 2024); multiple environments only (Van Der Vegt, Jansen, and Junges 2023; Chatterjee et al. 2025); or in online learning (Liu et al. 2022; Chen et al. 2023). Our work focuses on such a subclass, namely revealing POMDPs, which has been studied in the literature.
Avrachenkov, Dhiman, and Kavitha (2025) studied strongly-connected revealing POMDPs and restricting attention to the set of belief-stationary policies. They proved that there is an optimal policy within this set, whose induced belief dynamics are contracting and reach a positive recurrent class of beliefs in finite time. They also provided an infinite-dimensional linear program that characterizes the optimal belief-stationary policy. They considered an application of strongly-connected revealing blind MDPs. (Chen and Liew 2023) introduced the model of revealing blind MDPs. They consider the discounted objective and present algorithms based on finite-horizon approximation to compute the discounted value. Our work addresses the open computational analysis questions for revealing POMDPs.
2 Preliminaries
Notation
For a positive integer $n$ the set $\{1,2,\ldots,n\}$ is denoted by $[n]$. Sets and correspondences are denoted by calligraphic letters, e.g., $\mathcal{S},\mathcal{A},\mathcal{Z}$. Elements of these sets are denoted by lowercase letters, e.g., $s,a,z$. Random elements with values in these sets are denoted by uppercase letters, e.g., $S,A,Z$. The set of probability measures over a finite set $\mathcal{S}$ is denoted by $\Delta(\mathcal{S})$. The Dirac measure at some element $s\in\mathcal{S}$ is denoted by $\mathds{1}[s]$. The support of a probability measure $b\in\Delta(\mathcal{S})$ is denoted by $\operatorname{supp}(b)$.
Definition 1 (POMDP).
A POMDP is a tuple $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, where:
  • $\mathcal{S}$ is a finite set of states;
  • $\mathcal{A}$ is a finite set of actions;
  • $\mathcal{Z}$ is a finite set of signals;
  • $\delta\colon\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S}\times\mathcal{Z})$ is a probabilistic transition function that, given a state $s$ and an action $a$, returns the probability distribution over the successor states and signal;
  • $b_{0}\in\Delta(\mathcal{S})$ is the initial belief over the states.
Markov Decision Processes (MDPs) are POMDPs in which the signal corresponds to the state (Puterman 2014). Formally, an MDP is denoted by $M=(\mathcal{S},\mathcal{A},\delta,b_{0})$ with transition function $\delta\colon\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$.
Dynamic
Given a POMDP, a controller knows all defining parameters. At the beginning, nature draws a state $S_{0}\sim b_{0}$, which is not informed to the controller. Then, at each step $t\geq 0$, the controller chooses an action $A_{t}\in\mathcal{A}$, possibly at random. In response, nature draws the next state and signal $(S_{t+1},Z_{t+1})\sim\delta(S_{t},A_{t})$. The signal $Z_{t+1}$ is revealed to the controller, but the state $S_{t+1}$ is not announced. A play (or a path) in the POMDP is an infinite sequence $\rho=(s_{0},a_{0},z_{1},s_{1},a_{1},z_{2},s_{2},a_{2},\ldots)$ of states, actions, and signals such that, for all $t\geq 0$, we have $\delta(s_{t},a_{t})(s_{t+1},z_{t+1})>0$. The set of all plays is denoted by $\Omega$.
Belief
At each step $t\geq 0$, the controller’s (random) belief about the current state can be computed using Bayes’ rule and is denoted by $B_{t}\in\Delta(\mathcal{S})$. In the cases where the controller knows the exact state the controller’s belief is a Dirac measure $\mathds{1}[s]$ at some state $s\in\mathcal{S}$.
Definition 2 (Revealing POMDP).
A POMDP is revealing if, each time a state is visited, the state is also announced to the controller with positive probability. Formally, for each state there is a designated signal, i.e., $\mathcal{S}\subseteq\mathcal{Z}$, and, for all states $s,s^{\prime}\in\mathcal{S}$ and actions $a\in\mathcal{A}$,
$\sum_{z\in\mathcal{Z}}\delta(s,a)(s^{\prime},z)>0$
$\qquad\implies\sum_{\widetilde{s}\in\mathcal{S}}\delta(s,a)(\widetilde{s},s^{\prime})=\delta(s,a)(s^{\prime},s^{\prime})>0\,.$
Revealing POMDPs coincide with the class of strongly revealing defined in (Belly et al. 2025). Indeed, if the controller observes a signal $s\in\mathcal{Z}$, then their next belief is $\mathds{1}[s]$, which corresponds to a revelation of the state.
Remark 1 (Signals).
The signaling structure of POMDPs is modeled in different ways in the literature. We comment on two general cases.
  • The controller may receive a set of signals at each step with a transition function $\delta\colon\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S}\times 2^{\mathcal{Z}})$. In that case, we consider each set of signals as one signal $\widetilde{\mathcal{Z}}\coloneqq 2^{\mathcal{Z}}$ and reduce to our model. Even with exponentially many signals encoded in a polynomial size input, our EXPTIME upper bound holds.
  • In general POMDPs, it is enough to consider transition functions that pair each state with only one signal, known as deterministic observation, see (Chatterjee, Chmelík, and Tracol 2016, Remark 4). For revealing POMDPs, deterministic observations do not capture the general case because they correspond to MDPs.
Policies
A policy defines how a controller selects actions based on all the information available up to a given step. Formally, a (history-dependent randomized) policy is a function $\sigma\colon(\mathcal{A}\times\mathcal{Z})^{*}\to\Delta(\mathcal{A})$. The set of all policies is denoted by $\Sigma$. A policy is pure if it prescribes deterministic actions, i.e., it corresponds to a function $\sigma\colon(\mathcal{A}\times\mathcal{Z})^{*}\to\mathcal{A}$.
Probability Measures
For a finite prefix of a play, an element in $(\mathcal{S}\times\mathcal{A})^{*}$, its cone is the set of plays with it as their prefix. Given a policy $\sigma$ and an initial belief $b_{0}$, the unique probability measure over Borel sets of infinite plays obtained given $\sigma$ is denoted by $\mathbb{P}^{\sigma}_{b_{0}}(\cdot)$, which is defined by Carathéodory’s extension theorem by extending the natural definition over cones of plays (Billingsley 2012).
Objectives
An objective in a POMDP is a Borel set of plays $\Phi\subseteq\Omega$ in the Cantor topology on $\Omega$ (Kechris 1995). We consider objectives in the first $2\nicefrac{{1}}{{2}}$ levels of the Borel hierarchy, including the parity objective which expresses all $\omega$-regular objectives (Thomas 1997). Denote the state at time $t$ by $s_{t}$ and set of states that occur infinitely often in a play $\rho$ by $\mathcal{I}(\rho)\coloneqq\{s\in\mathcal{S}:\forall t\geq 0\;\exists\widetilde{t}\geq t\quad s=s_{\widetilde{t}}\in\rho\}$. We consider the following objectives.
  • Reachability: Given a set $\mathcal{X}\subseteq\mathcal{S}$ of target states, the reachability objective requires that a target state is visited at least once. Formally, the reachability objective is $\operatorname{Reach}(\mathcal{X})\coloneqq\{\rho\in\Omega:\exists t\geq 0\quad s_{t}\in\mathcal{X}\}$.
  • Parity: Given a $d\geq 0$ and a function $\operatorname{pri}\colon\mathcal{S}\to\{0,1,\ldots,d\}$ of priorities, the parity objective requires that the smallest priority that appears infinitely often is even. Formally, $\operatorname{Parity}\coloneqq\left\{\rho\in\Omega:\min\{\operatorname{pri}(s):s\in\mathcal{I}(\rho)\}\text{ is even}\right\}$.
Value
The value of an objective in a POMDP is the maximum probability a controller can guarantee to satisfy the objective. Formally, given an objective $\Phi\subseteq\Omega$, the value is a function of the initial belief $\operatorname{val}_{\Phi}\colon\Delta(\mathcal{S})\to[0,1]$ defined by $\operatorname{val}_{\Phi}(b)\coloneqq\sup_{\sigma\in\Sigma}\mathbb{P}^{\sigma}_{b}(\rho\in\Phi)$. Denote the reachability and parity values by $\operatorname{val}_{\textnormal{R}(\mathcal{X})}$ and $\operatorname{val}_{\textnormal{P}}$, respectively. We may omit $\mathcal{X}$ in $\operatorname{val}_{\textnormal{R}(\mathcal{X})}$ if it is clear from the context.
Approximately Optimal Policies
Given $\varepsilon\geq 0$ and an objective $\Phi\subseteq\Omega$, a policy $\sigma\in\Sigma$ is $\varepsilon$-optimal if it guarantees the value up to an additive error $\varepsilon$, i.e., if $\mathbb{P}_{b}^{\sigma}(\rho\in\Phi)\geq\operatorname{val}_{\Phi}(b)-\varepsilon$. In particular, we call 0-optimal policies simply optimal.
Computational Analysis
The computational analysis problems for POMDPs with an objective $\Phi$ are:
  • Almost-sure analysis asks whether the objective can be satisfied with probability $1$, i.e., $\exists\sigma\in\Sigma$ such that $\mathbb{P}_{b_{0}}^{\sigma}(\rho\in\Phi)=1$.
  • Limit-sure analysis asks whether the objective can be satisfied with probability arbitrarily close to 1, i.e., $\forall\varepsilon>0$ $\exists\sigma\in\Sigma$ such that $\mathbb{P}_{b_{0}}^{\sigma}(\rho\in\Phi)\geq 1-\varepsilon$ or equivalently whether $\operatorname{val}_{\Phi}(b_{0})=1$.
  • Quantitative analysis asks to compute an approximation of the optimal value, i.e., for all $\varepsilon>0$, provide $v\in[0,1]$ such that $\left|v-\operatorname{val}_{\Phi}(b_{0})\right|\leq\varepsilon$.
3 Overview
We present an overview of our approach and the results.
3.1 Overview of Approach
To solve the qualitative and quantitative analysis of parity objectives, we introduce a new objective called belief-reachability and show that the parity value coincides with the belief-reachability value to the set of almost-sure winning parity states.
Belief-Reachability
Given a set of target states $\mathcal{X}\subseteq\mathcal{S}$, the belief-reachability objective requires that a target state is visited and the controller has complete knowledge of the state at that step at least once. Formally, the set of Dirac beliefs on $\mathcal{X}\subseteq\mathcal{S}$ is denoted by $\mathcal{D}_{\mathcal{X}}\coloneqq\left\{b\in\Delta(\mathcal{S}):\exists s\in\mathcal{X}\textnormal{ such that }b=\mathds{1}[s]\right\}$. Then, the belief-reachability objective is $\operatorname{Belief-Reach}(\mathcal{X})\coloneqq\{\rho\in\Omega:\exists t\geq 0\quad B_{t}(\rho)\in\mathcal{D}_{\mathcal{X}}\}$. Denote the belief-reachability value by $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}$. We may omit $\mathcal{X}$ in $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}$ if it is clear from the context.
Remark 2 (Generality of belief-reachability).
In general POMDPs with reachability objectives, without loss of generality, target states can be considered absorbing, i.e., the dynamic remains in the same state once a target state is visited. Furthermore, one can reduce to the case where there is only one target state. These simplifications affect the dynamic, but neither optimal policies nor the reachability value. Belief-reachability generalizes reachability objectives as follows. Given a POMDP $P$ with an absorbing target state $s^{*}$, consider a copy of $P$ but add a new action and two absorbing states $\top$ and $\bot$. After playing the new action, the state moves to $\top$ if the state was in $s^{*}$, and to $\bot$ otherwise. Moreover, the controller is announced which of these states is reached. Then, the belief-reachability value to $\top$ coincides with the original reachability value. Indeed, for an approximately optimal policy of $P$, play the new action after sufficiently many steps to obtain approximately the same value in the belief-reachability objective.
3.2 Overview of Results
Results for Belief-Reachability Objectives
Our main results for belief-reachability objectives are the following, which are proved in Section 4.
Theorem 1.
Quantitative analysis for belief-reachability objectives for revealing POMDPs is in EXPTIME.
By Remark 2, reachability objectives reduce to belief-reachability objectives. Therefore, they have the following consequence.
Corollary 2.
Quantitative analysis for reachability objectives for revealing POMDPs is in EXPTIME.
Results for Parity Objectives
Our main results for parity objectives are the following, which are proved in Section 5.
Theorem 3.
Quantitative analysis for parity objectives for revealing POMDPs is in EXPTIME.
Theorem 3 generalizes Corollary 2 to parity objectives, and follows from a reduction of parity objectives to belief-reachability to a set that can be computed in EXPTIME, see Lemma 15.
Theorem 4.
Optimal policies exist for parity objectives for revealing POMDPs.
The following results follow from (Belly et al. 2025, Theorem 3) and Theorem 4.
Corollary 5.
For revealing POMDPs with parity objectives, limit-sure and almost-sure winning coincide, and limit-sure analysis is in EXPTIME-complete.
4 Belief-Reachability Objectives
In this section, we prove Theorem 1, i.e., that the quantitative analysis for belief-reachability objectives is in EXPTIME. We proceed in five steps:
  • We introduce reliable actions, which preserve the current belief-reachability value. Proposition 6 proves that every belief has a reliable action.
  • Lemma 7 shows that stopping policies that play only reliable actions are optimal.
  • Lemma 8 shows that playing reliable actions uniformly at random is stopping.
  • Lemma 10 upper bounds the horizon needed to approximate the belief-reachability value, using stopping optimal policies.
  • Lemma 11 presents a point-based dynamic programming algorithm to approximate the finite-horizon belief-reachability value.
Definition 3 (Reliable Action).
Consider a POMDP with belief-reachability objectives. An action $a\in\mathcal{A}$ is reliable if it preserves the belief-reachability value. Formally, for each belief $b\in\Delta(\mathcal{S})$, define the set of reliable actions $\mathcal{R}(b)$ by
$\mathcal{R}(b)\coloneqq\left\{a\in\mathcal{A}:\mathbb{E}^{a}_{b}(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{1}))=\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b)\right\}\,.$
Because the set of actions $\mathcal{A}$ is finite, we have the following result.
Proposition 6.
The set of reliable actions is nonempty.
Definition 4 (Terminal State).
Consider a POMDP and a set of target states $\mathcal{X}\subseteq\mathcal{S}$. A state $s\in\mathcal{S}$ is called terminal (for $\mathcal{X}$) if $s\in\mathcal{X}$ or if $\mathcal{X}$ cannot be reached from $s$ with positive probability, i.e., $\sup_{\sigma\in\Sigma}\mathbb{P}_{\mathds{1}[s]}^{\sigma}(\exists t\geq 0\quad\operatorname{supp}(B_{t})\cap\mathcal{X}\not=\emptyset)=0$. The set of terminal states is denoted by $\mathcal{T}$.
Definition 5 (Stopping Policy).
Consider a POMDP, a set of target states $\mathcal{X}\subseteq\mathcal{S}$, and a corresponding set of terminal states $\mathcal{T}\subseteq\mathcal{S}$. For $n\in\mathbb{N}$ and $q>0$, a policy $\sigma\in\Sigma$ is called $(n,q)$-stopping (for $\mathcal{X}$) if within $n$ steps the belief collapses to a Dirac mass on a terminal state with probability at least $q$. Formally, from every initial belief $b\in\Delta(\mathcal{S})$,
$\mathbb{P}_{b}^{\sigma}(\exists t\leq n\quad B_{t}\in\mathcal{D}_{\mathcal{T}})\geq q\,.$
We say that a policy is stopping if it is stopping for some $n$ and $q>0$.
Similar to (Andersson and Miltersen 2009, Lemma 5) that considers finite duration games, it is easy to see that, if a policy is stopping and uses only reliable actions, then it is optimal.
Lemma 7.
Consider a POMDP with belief-reachability objectives. If a policy is stopping and uses only reliable actions, then it is optimal.
Proof Sketch.
Consider a stopping policy $\sigma\in\Sigma$ which uses only reliable actions. Because every action chosen by $\sigma$ is reliable, the belief‑reachability values $\operatorname{val}_{\textnormal{BR}}(B_{t})$ forms a martingale, i.e., $\operatorname{val}_{\textnormal{BR}}(b_{0})=\mathbb{E}_{b_{0}}^{\sigma}\bigl(\operatorname{val}_{\textnormal{BR}}(B_{t})\bigr)$. Moreover, since the policy $\sigma$ is stopping, it reaches a Dirac belief on a terminal state in finite time with probability 1 and the hitting time $\tau$ of $\mathcal{D}_{\mathcal{T}}$ is almost-surely finite. Therefore, we have $\operatorname{val}_{\textnormal{BR}}(b_{0})=\mathbb{E}_{b_{0}}^{\sigma}\bigl(\operatorname{val}_{\textnormal{BR}}(B_{\tau})\bigr)$. ∎
Minimal Positive Transition
Denote $\delta_{\min}$ the minimum non-zero probability in the transition function, i.e.,
$\delta_{\min}\coloneqq\min\Bigl\{\delta(s,a)(s^{\prime},z)\;:\;\forall s,\,s^{\prime}\in\mathcal{S},\;a\in\mathcal{A},\;z\in\mathcal{Z}$
$\delta(s,a)(s^{\prime},z)>0\Bigr\}.$
Lemma 8.
Consider a revealing POMDP with belief-reachability objectives. The policy $\sigma$ that plays reliable actions uniformly at random is $(n,q)$-stopping with parameters $n\coloneqq|\mathcal{S}|+2$ and $q\coloneqq\delta_{\min}^{2}\left(\delta_{\min}/|\mathcal{A}|\right)^{|\mathcal{S}|}$.
Proof Sketch.
By Proposition 6, every belief has a reliable action, so the policy $\sigma$ is well-defined. Build the layered sets $\mathcal{S}_{0}\coloneqq T$, $\mathcal{S}_{t+1}\coloneqq\{s:\exists a\in\mathcal{R}(\mathds{1}[s])\text{ with a positive‑probability move to }\mathcal{S}_{t}\}$. The sequence covers all states after at most $|\mathcal{S}|$ iterations, i.e., $\mathcal{S}_{|\mathcal{S}|}=\mathcal{S}$, because, if there were states outside, then they require to use unreliable actions to get to terminal states, which lowers their value and forms a contradiction. Under the policy $\sigma$ that plays every reliable action uniformly, (a) the first “reveal” of the Dirac belief happens with probability $\delta_{\min}$; (b) each step toward the next layer then succeeds with probability at least $\delta_{\min}/|\mathcal{A}|$; and (c) the final “reveal” of the Dirac belief happens with probability $\delta_{\min}$. Hence, within $n\coloneqq|\mathcal{S}|+2$ steps, a Dirac belief is reached on a terminal state with probability $q\coloneqq\delta_{\min}^{2}\left(\delta_{\min}/|\mathcal{A}|\right)^{|\mathcal{S}|}$, which proves that $\sigma$ is $(n,q)$-stopping. ∎
We deduce directly from Lemma 7 and Lemma 8 the next result.
Corollary 9.
Every revealing POMDP with belief-reachability objectives has an optimal policy.
We turn to the quantitative analysis for the belief-reachability value. Note that the existence of optimal stopping policies implies that this value can be approximated using a finite horizon.
Lemma 10.
Consider a POMDP with belief-reachability objectives and an $(n,q)$-stopping optimal policy $\sigma$. For all $\varepsilon>0$, the finite horizon $T\coloneqq n\left\lceil\frac{\log(\varepsilon)}{\log(1-q)}\right\rceil$ is such that
$\mathbb{P}^{\sigma}_{b_{0}}(\exists t\leq T\quad B_{t}\in\mathcal{D}_{\mathcal{T}})\geq\operatorname{val}_{\textnormal{BR}}(b_{0})-\varepsilon\,.$
Proof Sketch.
Split the play into blocks of $n$ steps. Because the optimal policy $\sigma$ is $(n,q)$-stopping, each block reaches a Dirac belief on a terminal state with probability at least $q$. Thus for the hitting time $\tau$ of $\mathcal{D}_{\mathcal{T}}$ can be bounded by $\mathbb{P}_{b_{0}}^{\sigma}(\tau>kn)\leq(1-q)^{k}$. Therefore, $\tau$ has a geometric tail and is almost surely finite. Choose $T=n\left\lceil\dfrac{\log(\varepsilon)}{\log(1-q)}\right\rceil$ so that $\mathbb{P}^{\sigma}_{b_{0}}(\tau>T)\leq\varepsilon$, showing that a horizon of $T$ steps suffices to approximate the belief-reachability value. ∎
Reduction to Finite Horizon
To solve the quantitative analysis for revealing POMDPs with belief-reachability objectives, it suffices to compute the finite horizon value for a horizon that is exponentially large on the input size. Indeed, by Lemma 8, there exists an $(n,q)$-stopping optimal policy with $n=|\mathcal{S}|+2$ and $q=(\delta_{\min})^{2}(\delta_{\min}/|\mathcal{A}|)^{|\mathcal{S}|}$. Hence, by Lemma 10, for every $\varepsilon>0$, we can consider $T=n\left\lceil\frac{\log(\varepsilon)}{\log(1-q)}\right\rceil=O\left(|\mathcal{S}||\mathcal{A}|^{|\mathcal{S}|}\log(1/\varepsilon)/\delta_{\min}^{|\mathcal{S}|+2}\right)$, which is at most exponentially large in the input size.
Previous Approaches to Finite Horizon
Finite horizon objectives have been studied by a long time. A naive approach to the quantitative analysis for belief-reachability objectives would take the time bound $T$ in Lemma 10 and solve a (fully observable) MDP with $|\mathcal{A}\times\mathcal{S}|^{T}$ states. This approach implies a 2EXPTIME complexity upper bound. A fundamental result states that, for horizons that are polynomially large with respect to the input, computing the finite horizon value of POMDPs is PSPACE-complete (Papadimitriou and Tsitsiklis 1987, Theorem 6, page 448). The technique, instead of listing explicitly all exponentially many histories, compactly represents them using nondeterminism and space proportional to the horizon. The conclusion follows from PSPACE being closed under nondeterminism by Savitch’s theorem. For exponentially large horizons, this technique leads to an EXPSPACE upper bound. Instead, we prove an EXPTIME upper bound.
Point-based algorithms were introduced for POMDPs by (Pineau, Gordon, and Thrun 2003) as an alternative to approximating the value of POMDPs by considering a fixed subset of beliefs and projected belief updates. For a survey on point-based algorithms see (Shani, Pineau, and Kaplow 2013; Walraven and Spaan 2019), which focuses on finite horizon objectives but does not provide the complexity upper bound we require.
Finite-Horizon Belief-Reachability
Let $T\in\mathbb{N}$ be a finite horizon. The $T$-step belief-reachability value to $\mathcal{X}\subseteq\mathcal{S}$ is defined by $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T}(b)\coloneqq\max_{\sigma\in\Sigma}\mathbb{P}_{b}^{\sigma}(\exists t\leq T\;\,B_{t}\in\mathcal{D}_{\mathcal{X}}).$
Lemma 11.
Approximating the $T$-step belief-reachability value of revealing POMDPs up to an additive error of $\varepsilon>0$ can be computed in exponential time, formally, in time $O\left(T^{|\mathcal{S}|}|\mathcal{S}|^{|\mathcal{S}|+1}|\mathcal{A}||\mathcal{Z}|/\varepsilon^{(|\mathcal{S}|-1)}\right)$.
Proof Sketch.
We approximate the $T$-step belief-reachability value by discretizing the belief simplex and applying projected Bellman updates over this grid. Define a $k$-uniform grid $G_{k}\subseteq\Delta(\mathcal{S})$ so that every belief $b\in\Delta(\mathcal{S})$ is at $L_{1}$-distance at most $\varepsilon/(T+1)$ from some grid point $\Pi_{k}(b)$, where $k=\left\lceil(T+1)|\mathcal{S}|/\varepsilon\right\rceil$. Then, compute the Bellman updates on grid points, carefully projecting back to $G_{k}$ after each update, for $T$ steps. By induction on the horizon, the error at each step accumulates by at most $\varepsilon/(T+1)$, yielding a total error at most $\varepsilon$. The grid has size $O((T|\mathcal{S}|/\varepsilon)^{|\mathcal{S}|-1})$, and each Bellman update takes $O(|\mathcal{S}|^{2}|\mathcal{A}||\mathcal{Z}|)$ time, so the total running time is $O\left(T^{|\mathcal{S}|}|\mathcal{S}|^{|\mathcal{S}|+1}|\mathcal{A}||\mathcal{Z}|\varepsilon^{-(|\mathcal{S}|-1)}\right)$, which is exponential in the input size. ∎
The following result follows from Lemmas 10 and $1$. See 1
5 Parity Objectives
5.1 Reduction to Belief-Reachability Objectives
In the sequel, we show that, for revealing POMDPs, the parity objective value coincides with the belief-reachability to the set of Dirac on the almost‐sure winning parity states. This proof uses the standard notion of end components in MDPs, introduced in (Alfaro 1998); see also (Baier and Katoen 2008, Section 10.6.3).
Underlying MDPs of POMDPs
For a POMDP $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, we define its underlying MDP as $M=(\mathcal{S}\times(\mathcal{Z}\cup\{\square\}),\mathcal{A},\widetilde{\delta},\widetilde{b}_{0})$, where the transition function is defined as, for all $s,s^{\prime}\in\mathcal{S},z\in\mathcal{Z}\cup\{\square\},z^{\prime}\in\mathcal{Z}$,
$\widetilde{\delta}\bigl((s,z),a\bigr)\bigl((s^{\prime},z^{\prime})\bigr)\coloneqq\delta(s,a)(s^{\prime},z^{\prime})\,$$\qquad$
and the initial belief is defined as $\widetilde{b}_{0}\bigl((s,\square)\bigr)\coloneqq b_{0}(s)$ for all states $s\in\mathcal{S}$. This construction captures the entire dynamic of $P$. We use this notion to analyze the policies derived from the original POMDP.
End Components
Let $M=(\mathcal{S},\mathcal{A},\delta,b_{0})$ be an MDP. An end component is a pair $\mathcal{U}=(\mathcal{Q},\mathcal{E})$ where $\mathcal{Q}\subseteq\mathcal{S}$ is a subset of states and $\mathcal{E}\colon\mathcal{Q}\rightrightarrows\mathcal{A}$ assigns each state to a set of nonempty actions such that
  • Closedness. For all states $q\in\mathcal{Q}$ and actions $a\in\mathcal{E}(q)$, we have $\operatorname{supp}(\delta(q,a))\subseteq\mathcal{Q}$; and
  • Strong connectivity. The directed graph with states $\mathcal{Q}$ and edges $(q,q^{\prime})$ where there exists $a\in\mathcal{E}(q)$ such that $\delta(q,a)(q^{\prime})>0$ is strongly connected.
A play $\rho$ visits infinitely often an end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})$, denoted by $\mathcal{I}(\rho)=(\mathcal{Q},\mathcal{E})$, if
  • The set of states visited infinitely often in $\rho$ is $\mathcal{Q}$; and
  • For all state $q\in\mathcal{Q}$, the set of actions played infinitely often when in $q$ is $\mathcal{E}(q)$.
[…]
The following statement is a fundamental result of end components in MDPs.
Lemma 12 ((Alfaro 1998, Theorems 3.1 and 3.2)).
For MDPs, the following assertions hold.
  • For every end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})$ and state $q\in\mathcal{Q}$, there exists a policy $\sigma$ such that
    $\mathbb{P}_{\mathds{1}[q]}^{\sigma}\left(\mathcal{I}(\rho)=\mathcal{U}\right)=1\,.$
  • For all policies $\sigma$, we have
    $\mathbb{P}_{b_{0}}^{\sigma}\left(\mathcal{I}(\rho)\textnormal{ is an end component}\right)=1\,.$
The following result relates end components in MDPs to policies in general POMDPs.
Lemma 13.
Consider a POMDP $P$ and its underlying MDP $M$. For every reachable end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})$ of $M$, i.e., there exists a policy $\sigma$ on $P$ such that $\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})>0$, we have that, for all states $q\in\mathcal{Q}$, actions $a\in\mathcal{E}(q)$, and signals $z\in\mathcal{Z}$, there exists a policy $\sigma_{q,a,z}$ on $P$ such that
$\mathbb{P}_{b}^{\sigma_{q,a,z}}(\mathcal{I}(\rho)\subseteq\mathcal{U})\;=\;1\,$,
where $b$ is the belief after starting with belief $\mathds{1}[q]$, playing action $a$, and receiving signal $z$.
Proof Sketch.
Consider a reachable end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})$ of $M$ with corresponding policy $\sigma$, i.e., $\mathbb{P}^{\sigma}_{b_{0}}(\mathcal{I}(\rho)=\mathcal{U})>0$. On the event $\{\mathcal{I}(\rho)=\mathcal{U}\}$, every state-action pair $(q,a)$ with $q\in\mathcal{Q}$ and $a\in\mathcal{E}(q)$ is taken infinitely often, and each such occurrence results in at least one signal $z$ that occurs with positive probability. Fix $q\in\mathcal{Q}$, $a\in\mathcal{E}(q)$ and a signal $z$ that can follow $(q,a)$. By contradiction, assume that from the belief $b$ obtained after starting with $\mathds{1}[q]$, playing action $a$, and receiving signal $z$, no policy can stay inside $\mathcal{U}$ almost surely. Then, starting from $b$, any policy has a positive, state–independent strictly positive lower bound on the probability of leaving $\mathcal{Q}$ within a bounded number of steps. Because $(q,a)$ is visited infinitely often on the event $\{\mathcal{I}(\rho)=\mathcal{U}\}$, these lower‑bounded leaving probabilities accumulate, driving the probability of ever leaving $\mathcal{Q}$ to $1$. This contradicts $\mathbb{P}^{\sigma}_{b_{0}}(\mathcal{I}(\rho)=\mathcal{U})>0$. Hence, such a leaving bound cannot exist: there must be a policy $\sigma_{q,a,z}$ that, from $b$, keeps the play inside $\mathcal{Q}$ (and therefore inside $\mathcal{U}$) with probability $1$. ∎
The following example demonstrates that the above result is tight in the sense that it can not guarantee a policy $\sigma_{q,a,z}$ such that $\mathbb{P}_{b}^{\sigma_{q,a,z}}\bigl(\mathcal{I}(\rho)=\mathcal{U}\bigr)=1$.
Example 1.
Consider the example presented in (Chatterjee, Saona, and Ziliotto 2021, Example 4.3) of a POMDP with only one possible signal in Figure 1, commonly known as a blind MDP. It has four states $\mathcal{S}=\{s_{0},s_{1},\bot,\top\}$ and two actions: wait ($w$) and commit ($c$), i.e., $\mathcal{A}=\{w,c\}$. The priority function is defined as
$\operatorname{pri}(s_{0})\coloneqq 1,\;\operatorname{pri}(s_{1})\coloneqq 1,\;\operatorname{pri}(\bot)\coloneqq 1,\;\operatorname{pri}(\top)=0\,.$
The transitions are defined as follows. (a) Under action $w$, state $s_{0}$ moves either to $s_{1}$ or loops, both with probability $1/2$; states $s_{1}$ and $\bot$ loop; and state $\top$ moves to $s_{0}$. (b) Under action $c$, state $s_{0}$ moves to $\bot$; state $s_{1}$ moves $\top$; state $\bot$ loops; and state $\top$ moves to $s_{0}$. The only end component of the underlying MDP that satisfies the parity condition is
$\mathcal{U}=\bigl(\{s_{1},\top\},\;\{(s_{0},w),(s_{1},w),(s_{1},c),(\top,w),(\top,c)\}\bigr)\,.$
Note that, for every $\varepsilon$‑optimal policy $\sigma$, we have that $\mathbb{P}_{\mathds{1}[s_{0}]}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})\geq 1-\varepsilon$. This is achieved only by policies that wait long enough before committing, which requires unbounded memory. Conversely, no policy can guarantee $\mathbb{P}_{\mathds{1}[s_{0}]}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})=1$. Indeed, whenever action $c$ is taken, the belief on state $s_{0}$ is positive and therefore the state absorbs at $\bot$ with positive probability. The best that can be achieved almost surely is to stay forever in the end component $\widetilde{\mathcal{U}}=\bigl(\{s_{1}\},\{(s_{1},w)\}\bigr),$ whose minimal priority is 1 and therefore violates the parity condition. ∎
$s_{0}$$s_{1}\bot\topwwwwwwccccw$,cw,cw,cw,c
Figure 1: Example of a classic general POMDP that requires infinite memory policies for parity objectives. Edges represent a positive probability transition between states when the corresponding action in its label is used.
We now show that in revealing POMDPs, the above result can be extended to guarantee a policy $\sigma_{q}$ such that $\mathbb{P}_{\mathds{1}[q]}^{\sigma_{q}}\bigl(\mathcal{I}(\rho)=\mathcal{U}\bigr)=1\,.$
Lemma 14.
Consider a revealing POMDP $P$ and its underlying MDP $M$. For every policy $\sigma$ on $P$ and end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})$ of $M$ such that $\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})>0$, we have, for all states $q\in\mathcal{Q}$, there exists a policy $\sigma_{q}$ on $P$ such that
$\mathbb{P}_{\mathds{1}[q]}^{\sigma_{q}}(\mathcal{I}(\rho)=\mathcal{U})\;=\;1\,.$
Proof Sketch.
[…]
We are ready to present the reduction of parity to belief-reachability of the almost-sure winning states for parity.
Lemma 15.
Consider a revealing POMDP with parity objectives. Denote the set of states for which, if the initial belief were a Dirac on that state, then the parity condition can be satisfied almost-surely by $\mathcal{X}\coloneqq\left\{s\in\mathcal{S}:\exists\sigma\in\Sigma\quad\mathbb{P}_{\mathds{1}[s]}^{\sigma}(\operatorname{Parity})=1\right\}$. Then, the parity value coincides with the belief-reachability value to $\mathcal{X}$, i.e., for all beliefs $b\in\Delta(\mathcal{S})$,
$\operatorname{val}_{\textnormal{P}}(b)=\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b)\,.$
Proof sketch.
Consider a policy $\sigma\in\Sigma$. Consider its end components in the underlying MDP on $\mathcal{S}\times\mathcal{Z}$. Note that each end component either does satisfy or does not satisfy the parity condition. Moreover, if they satisfy the parity condition, then $\sigma$ is an almost-sure winning policy starting from any state inside the end component. Therefore, the probability of satisfying the parity condition under $\sigma$ corresponds to the belief-reachability to the end components where the parity condition is satisfied. In particular, in these end components, parity condition is satisfied almost-surely. We conclude because the policy is arbitrary. ∎
5.2 Proofs of Main Results
Building on Lemma 15, which reduces parity to belief‑reachability, and the EXPTIME procedure in Theorem 1 for approximating the belief-reachability value, we now (a) obtain an EXPTIME algorithm for approximating parity values (Theorem 3); and (b) show that limit‑sure and almost‑sure winning coincide (Theorem 4).
See 3
Proof Sketch.
By (Belly et al. 2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is EXPTIME. Therefore, this result is a direct implication of Lemma 15 and Theorem 1. ∎
See 4
Proof Sketch.
It is a direct implication of Corollary 9 and Lemma 15. ∎
Concluding Remarks
In this work, we consider revealing POMDPs which have been studied in the literature and provide decidability results for the fundamental computational problems for this model. Interesting directions for future work include exploring the practical applicability of our algorithms and extending our decidability results to other classes of POMDPs.
Acknowledgements
This work was supported by the ANRT under the French CIFRE Ph.D program in collaboration between NyxAir and Paris-Dauphine University (Contract: CIFRE N° 2022/0513), by the French Agence Nationale de la Recherche (ANR) under reference ANR-21-CE40-0020 (CONVERGENCE project), by Austrian Science Fund (FWF) 10.55776/COE12, and by the ERC CoG 863818 (ForM-SMArt) grant.
Appendix A Proofs of Section 4
See 6
Proof.
Consider a POMDP with initial belief $b_{0}\in\Delta(\mathcal{S})$ and set of target states $\mathcal{X}\subseteq\mathcal{S}$. By definition of the value, we have that $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b_{0})=\max_{a\in\mathcal{A}}\mathbb{E}^{a}_{b_{0}}(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{1}))$. Because the set of actions $\mathcal{A}$ is finite, there exists $a\in\mathcal{A}$ that achieves the maximum, which is a reliable action. ∎
See 7
Proof.
Consider a POMDP with initial belief $b_{0}\in\Delta(\mathcal{S})$ and belief-reachability objectives. Denote the set of target states and the corresponding set of terminal states by $\mathcal{X}\subseteq\mathcal{S}$ and $\mathcal{T}\subseteq\mathcal{S}$, respectively. Consider a policy $\sigma\in\Sigma$ that is stopping and uses only reliable actions. We show that $\sigma$ is optimal.
Denote the hitting time of terminal Dirac beliefs by $\tau\coloneqq\inf\{t\geq 0:B_{t}\in\mathcal{D}_{\mathcal{T}}\}$. Because the policy $\sigma$ is $(n,q)$-stopping, for every belief $b\in\Delta(\mathcal{S})$, we have that $\mathbb{P}_{b}^{\sigma}(\tau\leq n)\geq q$. Therefore, for every $k\geq 1$,
$\mathbb{P}_{b}^{\sigma}(\tau\leq kn)\geq 1-(1-q)^{k}\,.$
Taking the limit as $k$ goes to infinity, we deduce that $\mathbb{P}_{b}^{\sigma}(\tau<\infty)=1$.
Because $\sigma$ uses only reliable actions, we have that, for every $t\geq 1$,
$\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b)=\mathbb{E}^{\sigma}_{b}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{t})\right)\,.$
Therefore, for every $T\geq 1$ and $b\in\Delta(\mathcal{S})$,
$\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b)=\mathbb{E}^{\sigma}_{b}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{\tau})\mathds{1}_{\{\tau\leq T\}}\right)+\mathbb{E}^{\sigma}_{b}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{\tau})\mathds{1}_{\{\tau>T\}}\right)\,.$
Whenever $\tau\leq T$, we have that there exists $t\leq T$ such that $B_{t}\in\mathcal{D}_{\mathcal{T}}$, and thus $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{t})=\mathds{1}_{\{B_{t}\in\mathcal{D}_{\mathcal{X}}\}}$. Taking $T$ to infinity,
$\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b)$ $=\lim_{T\to\infty}\mathbb{E}^{\sigma}_{b}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{\tau})\mathds{1}_{\{\tau\leq T\}}\right)$
$=\lim_{T\to\infty}\mathbb{P}^{\sigma}_{b}(\exists t\leq T\quad B_{t}\in\mathcal{D}_{\mathcal{X}})$
$=\mathbb{P}^{\sigma}_{b}(\exists t\geq 0\quad B_{t}\in\mathcal{D}_{\mathcal{X}})\,.$
Therefore, the policy $\sigma$ is optimal. ∎
See 8
Proof.
Consider a revealing POMDP $P$ with belief-reachability objectives. Recall that, by Proposition 6, every belief has at least one reliable action, i.e., for every $b\in\Delta(\mathcal{S})$, we have that $\mathcal{R}(b)\neq\emptyset$. Let $\sigma$ be the policy that plays uniformly at random among all reliable actions. We show that $\sigma$ is $(n,q)$-stopping, i.e., for every initial belief $b\in\Delta(\mathcal{S})$,
$\mathbb{P}^{\sigma}_{b}(\exists t\leq n\quad B_{t}\in\mathcal{D}_{\mathcal{T}})\geq q\,$,
where
$n=|\mathcal{S}|+2\quad\textnormal{ and }\quad q=\delta_{\min}^{2}\left(\frac{\delta_{\min}}{|\mathcal{A}|}\right)^{|\mathcal{S}|}\,.$
Denote the set of terminal states by $\mathcal{T}\subseteq\mathcal{S}$. Define the set of states that can reach $\mathcal{T}$ using only reliable actions, indexed by the number of steps required, as follows. First, $\mathcal{S}_{0}\coloneqq\mathcal{T}$ and, for $t\geq 1$,
$\mathcal{S}_{t}\coloneqq\left\{s\in\mathcal{S}:\exists a\in\mathcal{R}\left(\mathds{1}[{s}]\right)\quad\sum_{\widetilde{s}\in\mathcal{S}_{t-1},z\in\mathcal{Z}}\delta(s,a)\left(\widetilde{s},z\right)>0\right\}\,.$
We claim that at most $|\mathcal{S}|$ steps are required to cover all states, i.e., $\mathcal{S}_{|\mathcal{S}|}=\mathcal{S}$.
Indeed, by contradiction, assume that the set of uncovered states $\mathcal{L}\coloneqq\mathcal{S}\setminus\mathcal{S}_{|\mathcal{S}|}$ is nonempty. Note that $(\mathcal{S}_{t})_{t\geq 0}$ is an increasing and bounded sequence and therefore reaches a fixpoint equal to $\mathcal{S}_{|\mathcal{S}|}$. On the one hand, for all state $s\in\mathcal{L}$, under reliable actions for the belief $\mathds{1}[s]$, the state can not move from $\mathcal{L}$ to $\mathcal{S}_{|\mathcal{S}|}$. Formally, for all $s\in\mathcal{L},a\in\mathcal{R}(\mathds{1}[s])$,
$\operatorname{supp}(\delta(s,a))\subseteq\mathcal{L}\times\mathcal{Z}\,.$
On the other hand, using unreliable actions imply decreasing the future value, i.e., for all $s\in\mathcal{L}$,
$\max_{a\in\mathcal{A}\setminus\mathcal{R}(\mathds{1}[s])}\mathbb{E}_{\mathds{1}[s]}^{a}(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{1}))<\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(\mathds{1}[s])\,.$
We show that there is a state in $\mathcal{L}$ with no $\varepsilon$-optimal policy for $\varepsilon$ small enough, which is a contradiction.
Consider the state $s\in\mathcal{L}$ with the highest belief-reachability value. Note that states with value zero are terminal, so $s$ has strictly positive value. Denote the minimum gap obtained by using unreliable actions by
$\zeta\coloneqq\min_{\begin{subarray}{c}s\in\mathcal{L},\\ a\notin\mathcal{R}(\mathds{1}[s])\end{subarray}}\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(\mathds{1}[s])-\mathbb{E}_{\mathds{1}[s]}^{a}(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{1}))>0\,.$
Take $\varepsilon\in(0,\min\{\zeta,\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(\mathds{1}[s])\})$ and consider a policy $\sigma_{\varepsilon}$ that is $\varepsilon$-optimal policy for the initial belief $\mathds{1}[s]$. Note that, because reliable actions do not connect $\mathcal{L}$ to $\mathcal{T}$ and $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(\mathds{1}[s])>0$, the policy $\sigma_{\varepsilon}$ must necessarily use unreliable actions. We show that $\sigma_{\varepsilon}$ is not $\varepsilon$-optimal, leading to a contradiction.
Denote the first time the action played has positive probability of leaving the set $\mathcal{L}$ by
$\tau\coloneqq\inf\left\{t\geq 0\;:\;\operatorname{supp}\bigl(\delta(S_{t},A_{t})\bigr)\notin\mathcal{L}\times\mathcal{Z}\right\}\,.$
Then, we bound the belief-reachability guaranteed by $\sigma_{\varepsilon}$ as follows.
$\mathbb{P}^{\sigma_{\varepsilon}}_{\mathds{1}[s]}(\exists t\in\mathbb{N}\quad B_{t}\in\mathcal{D}_{\mathcal{X}})$ $=\mathbb{E}_{\mathds{1}[s]}^{\sigma_{\varepsilon}}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(B_{\tau+1})\mathds{1}[\tau<\infty]\right)$ $(\mathcal{X}\cap\mathcal{L}=\emptyset)$
$\leq\mathbb{E}_{\mathds{1}[s]}^{\sigma_{\varepsilon}}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}\left(\mathds{1}[S_{\tau}]\right)\mathds{1}[\tau<\infty]\right)-\zeta$ $(\textnormal{Definition of $\zeta$})$
$\leq\operatorname{val}_{\textnormal{BR}(\mathcal{X})}\left(\mathds{1}[s]\right)-\zeta$ $(\textnormal{Choice of $s$})$
$<\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(\mathds{1}[s])-\varepsilon\,$, $(\textnormal{$\varepsilon<\zeta$})$
which contradicts the fact that $\sigma_{\varepsilon}$ is $\varepsilon$‑optimal. We conclude that $\mathcal{S}_{|\mathcal{S}|}=\mathcal{S}$.
We now prove that the policy $\sigma$ is stopping with parameters $n=|\mathcal{S}|+2$ and $q=\delta_{\min}^{2}\left(\frac{\delta_{\min}}{|\mathcal{A}|}\right)^{|\mathcal{S}|}$. Note that, because $P$ is revealing, for all $t\in[|\mathcal{S}|]$,
$s\in\mathcal{S}_{t}\quad\Rightarrow\quad\exists s^{\prime}\in\mathcal{S}_{t-1},a\in\mathcal{R}(\mathds{1}[s])\quad\delta(s,a)(s^{\prime},s^{\prime})>0\,.$
By recursion and because $\sigma$ plays all reliable actions uniformly at random, we conclude that, for all $s\in\mathcal{S}$,
$\mathbb{P}^{\sigma}_{\mathds{1}[s]}\left(\exists t\leq|\mathcal{S}|\quad S_{t}\in\mathcal{T}\right)\geq\left(\frac{\delta_{\min}}{|\mathcal{A}|}\right)^{|\mathcal{S}|}\,.$
and thus,
$\mathbb{P}^{\sigma}_{\mathds{1}[s]}\left(\exists t\leq|\mathcal{S}|+1\quad B_{t}\in\mathcal{D}_{\mathcal{T}}\right)\geq\delta_{\min}\left(\frac{\delta_{\min}}{|\mathcal{A}|}\right)^{|\mathcal{S}|}\,.$
To conclude, for all initial beliefs $b\in\Delta(\mathcal{S})$, after every action, the probability that the next belief is a Dirac at least $\delta_{\min}$. ∎
See 10
Proof.
Consider a POMDP with belief-reachability objectives and an $(n,q)$-stopping optimal policy $\sigma$. Denote the hitting time of the terminal set $\mathcal{T}\subseteq\mathcal{S}$ by
$\tau\coloneqq\inf\{t\geq 0:B_{t}\in\mathcal{D}_{\mathcal{T}}\}.$
Because $\sigma$ is $(n,q)$-stopping, we have that $\mathbb{P}^{\sigma}_{b_{0}}(\tau\leq n)\geq q$. Similarly, considering blocks of size $n$, for every integer $k\geq 1$, we have that $\mathbb{P}^{\sigma}_{b_{0}}(\tau>kn)\leq(1-q)^{k}$. Thus, $\tau$ is finite $\mathbb{P}^{\sigma}_{b_{0}}$-a.s.
For every $b_{0}\in\Delta(\mathcal{S})$, we have that
$\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b_{0})$ $=\mathbb{P}_{b_{0}}^{\sigma}(\exists t\geq 0\quad B_{t}\in\mathcal{D}_{\mathcal{X}})$ $\left(\textnormal{Definition of $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}$}\right)$
$=\mathbb{P}_{b_{0}}^{\sigma}(B_{\tau}\in\mathcal{D}_{\mathcal{X}})$ $(\tau<\infty)$
$=\mathbb{P}_{b_{0}}^{\sigma}(\tau\leq T,B_{\tau}\in\mathcal{D}_{\mathcal{X}})+\mathbb{P}_{b_{0}}^{\sigma}(\tau>T,B_{\tau}\in\mathcal{D}_{\mathcal{X}})$ $(\textnormal{Partitioning})$
$\leq\mathbb{P}_{b_{0}}^{\sigma}(\tau\leq T,B_{\tau}\in\mathcal{D}_{\mathcal{X}})+\mathbb{P}_{b_{0}}^{\sigma}(\tau>T)$ $(\textnormal{Inclusion})$
$\leq\mathbb{P}_{b_{0}}^{\sigma}(\exists t\leq T\quad B_{t}\in\mathcal{D}_{\mathcal{X}})+\mathbb{P}_{b_{0}}^{\sigma}(\tau>T)\,.$ $(\textnormal{Definition of $\tau$})$
We deduce that $\mathbb{P}^{\sigma}_{b_{0}}(\tau>T)\leq\varepsilon$ by taking $T=n\left\lceil\dfrac{\log(\varepsilon)}{\log(1-q)}\right\rceil$, which concludes the proof. ∎
See 11
Proof.
Consider a revealing POMDP with a set of target states $\mathcal{X}\subseteq\mathcal{S}$ and a finite horizon $T\in\mathbb{N}$. Denote the corresponding set of terminal states by $\mathcal{T}\subseteq\mathcal{S}$. Consider $b\in\Delta(\mathcal{S})$ and denote the L1-norm by $\|b\|_{1}\coloneqq\sum_{s\in\mathcal{S}}|b(s)|$.
Simplex Discretization: We first discretize the set of beliefs $\Delta(\mathcal{S})$. Let $k\geq 1$ and $\varepsilon>0$. Define the $k$-uniform grid of beliefs
$G_{k}\coloneqq\Bigl\{b\in\Delta(\mathcal{S}):$ $b(s)=\dfrac{n(s)}{k},n(s)\in\{0,\ldots,k\},\sum_{s=1}^{|\mathcal{S}|}n(s)=k\Bigr\}\,.$
Denote $\Pi_{k}\colon\Delta(\mathcal{S})\to G_{k}$ the projection function defined by
$\Pi_{k}(b)\coloneqq\operatorname*{arg\,min}_{\begin{subarray}{c}b^{\prime}\in G_{k}\\[2.0pt] \operatorname{supp}(b)=\operatorname{supp}(b^{\prime})\end{subarray}}\|b-b^{\prime}\|_{1}.$
By construction, it holds that, for every $b\in\Delta(\mathcal{S})$, there exists $\Pi_{k}(b)\in G_{k}$ with
$\|b-\Pi_{k}(b)\|_{1}\leq\dfrac{|\mathcal{S}|}{k},$
and taking $k=\left\lceil\dfrac{(T+1)|\mathcal{S}|}{\varepsilon}\right\rceil$, we have $\|b-\Pi_{k}(b)\|_{1}\leq\dfrac{\varepsilon}{T+1}$.
Bellman Equations: We now define the Bellman equations for the $T$-step belief-reachability value and then for the $T$-step belief-reachability value on the grid previously defined. For every $t\leq T$ and $b\in\Delta(\mathcal{S})$, define the bellman equation of the $t$-step belief reachability value in $P$ by
  • Base case: we have that
    $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{0}(b)\coloneqq\left\{\begin{array}[]{ll}1&\textnormal{ if }b=\mathds{1}[s]\textnormal{ for some }s\in\mathcal{X},\\ 0&\textnormal{ otherwise.}\end{array}\right$.
  • Induction case: we have that
    $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t}(b)$ $\coloneqq\mathds{1}[b\in\mathcal{D}_{\mathcal{X}}]+(1-\mathds{1}[b\in\mathcal{D}_{\mathcal{X}}])\max_{a\in\mathcal{A}}\sum_{z\in\mathcal{Z}}\sum_{s,s^{\prime}\in\mathcal{S}}b(s)\delta(s,a)(s^{\prime},z)\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}(\tau(b,a,z)),$
    where $\tau(b,a,z)(s^{\prime})\coloneqq\frac{\sum_{s\in\mathcal{S}}b(s)\delta(s,a)(s^{\prime},z)}{\sum_{s,s^{\prime}\in\mathcal{S}}b(s)\delta(s,a)(s^{\prime},z)}$.
We now define the value iteration that considers only beliefs on the grid $G_{k}$. More formally, for every $b\in\Delta(\mathcal{S})$, there exists $\Pi_{k}(b)\in G_{k}$ such that,
  • Base case: we have that
    $\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{0}(\Pi_{k}(b))\coloneqq\left\{\begin{array}[]{ll}1&\textnormal{ if }\Pi_{k}(b)=\mathds{1}[s]\textnormal{ for some }s\in\mathcal{X},\\ 0&\textnormal{ otherwise.}\end{array}\right$.
  • Induction case: we have that
    $\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t}(\Pi_{k}(b))\coloneqq\mathds{1}[\Pi_{k}(b)\in\mathcal{D}_{\mathcal{X}}]+(1-\mathds{1}[\Pi_{k}(b)\in\mathcal{D}_{\mathcal{X}}])\max_{a\in\mathcal{A}}\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t-1}(\Pi_{k}(B_{1}))\right).$
Lipschitz Property: We show that the $T$-step belief-reachability value is a Lipschitz function over beliefs with same support. For all beliefs $b,b^{\prime}\in\Delta(\mathcal{S})$ with $\operatorname{supp}(b)=\operatorname{supp}(b^{\prime})$, we have
$\left|\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T}(b)-\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T}(b^{\prime})\right|$
$\qquad=\left|\max_{a\in\mathcal{A}}\mathbb{E}_{b}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)-\max_{a\in\mathcal{A}}\mathbb{E}_{b^{\prime}}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)\right|$ $(\textnormal{Bellman optimality})$
$\qquad\leq\left|\max_{a\in\mathcal{A}}\left[\mathbb{E}_{b}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)-\mathbb{E}_{b^{\prime}}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)\right]\right|$ $(\text{Subadditivity of $\max$})$
$\qquad\leq\max_{a\in\mathcal{A}}\left|\mathbb{E}_{b}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)-\mathbb{E}_{b^{\prime}}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)\right|$ $(\text{Monotonicity of $\max$})$
$\qquad=\max_{a\in\mathcal{A}}\left|\sum_{s\in\mathcal{S}}\bigl(b(s)-b^{\prime}(s)\bigr)\mathbb{E}_{1[s]}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)\right|$ $\left(\textnormal{Affinity of }\mathbb{E}_{b}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)\right)$
$\qquad\leq\sum_{s\in\mathcal{S}}\left|b(s)-b^{\prime}(s)\right|=\left\|b-b^{\prime}\right\|_{1}\,$, $\left(0\leq\mathbb{E}_{1[s]}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{T-1}\left(B_{1}\right)\right)\leq 1\right)$
Error Bound: Take $k=\left\lceil\frac{(T+1)|\mathcal{S}|}{\varepsilon}\right\rceil$. Define for every $t\in\{0,\ldots,T\}$ and belief $b\in\Delta(\mathcal{S})$,
$E_{t}(b)\coloneqq\left|\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t}(b)-\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t}(\Pi_{k}(b))\right|$
We prove using an induction argument that, for every $t\leq T$ and $b\in\Delta(\mathcal{S})$,
$E_{t}(b)\leq\dfrac{(t+1)\varepsilon}{T+1}.$
Observe that when $t=0$, we have that $E_{0}(b)\leq\dfrac{\varepsilon}{T+1}$. Now, assume that $E_{t-1}\leq\frac{t\varepsilon}{T+1}$. Therefore, we have that, for every $b\in\Delta(\mathcal{S})$ and taking $k=\left\lceil\dfrac{(T+1)|\mathcal{S}|}{\varepsilon}\right\rceil$,
$E_{t}(b)$ $=\left|\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t}(b)-\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t}(\Pi_{k}(b))\right|$
$\leq\max_{a\in\mathcal{A}}\left|\mathbb{E}_{b}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}(B_{1})\right)-\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(\Pi_{k}(B_{1})\right)\right)\right|$
$\leq\max_{a\in\mathcal{A}}\left|\mathbb{E}_{b}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}(B_{1})\right)-\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(B_{1}\right)\right)\right|$
$\qquad+\max_{a\in\mathcal{A}}\left|\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(B_{1}\right)\right)-\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(\Pi_{k}(B_{1})\right)\right)\right|$
$\leq\left\|b-\Pi_{k}(b)\right\|_{1}$ (Lipschitz property)
$\qquad+\max_{a\in\mathcal{A}}\left|\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(B_{1}\right)\right)-\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(\Pi_{k}(B_{1})\right)\right)\right|$
$\leq\left\|b-\Pi_{k}(b)\right\|_{1}$
$\qquad+\max_{a\in\mathcal{A}}\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\left|\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(B_{1}\right)-\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(\Pi_{k}(B_{1})\right)\right|\right)$
$\leq\dfrac{\varepsilon}{T+1}+\max_{a\in\mathcal{A}}\mathbb{E}_{\Pi_{k}(b)}^{a}\left(\left|\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(B_{1}\right)-\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t-1}\left(\Pi_{k}(B_{1})\right)\right|\right)$ (Definition of $k$)
$\leq\dfrac{\varepsilon}{T+1}+\max_{a\in\mathcal{A}}\mathbb{E}_{b}^{a}(E_{t-1}(B_{1}))$ (Definition of $E_{t-1}$)
$\leq\dfrac{\varepsilon}{T+1}+\dfrac{t\varepsilon}{T+1}$ (Induction hypothesis)
$=\dfrac{(t+1)\varepsilon}{T+1},$
and the induction argument holds. Therefore, we obtain that $E_{T}(b)\leq\frac{(T+1)\varepsilon}{T+1}=\varepsilon$.
Complexity: For every action, the bellman update considers $O(|\mathcal{S}|^{2}|\mathcal{Z}|)$. Taking the maximum over actions, we get a complexity bound of $O(|\mathcal{S}|^{2}|\mathcal{A}||\mathcal{Z}|)$. We considered the update on a uniform grid $G_{k}\subseteq\Delta(\mathcal{S})$ with, for all $b\in\Delta(\mathcal{S})$, $\|b-\Pi_{k}(b)\|_{1}\leq\varepsilon/(T+1)$ by taking $k=\left\lceil\dfrac{(T+1)|\mathcal{S}|}{\varepsilon}\right\rceil$. The size of grid in the $(|\mathcal{S}|-1)$-simplex $\Delta(\mathcal{S})$ is thus
$\left|G_{k}\right|=\binom{k+|\mathcal{S}|-1}{|\mathcal{S}|-1}=O\left(\left(\tfrac{T|\mathcal{S}|}{\varepsilon}\right)^{|\mathcal{S}|-1}\right).$
Therefore, we deduce that
Total time complexity $=O\left((T+1)\cdot|G_{k}|\cdot|\mathcal{S}|^{2}||\mathcal{A}||\mathcal{Z}|\right)$
$=O\left((T+1)\cdot\left(\dfrac{T|\mathcal{S}|}{\varepsilon}\right)^{|\mathcal{S}|-1}\cdot|\mathcal{S}|^{2}||\mathcal{A}||\mathcal{Z}|\right)$
$=O\left(\dfrac{T^{|\mathcal{S}|}|\mathcal{S}|^{|\mathcal{S}|+1}|\mathcal{A}||\mathcal{Z}|}{\varepsilon^{|\mathcal{S}|-1}}\right),$
which completes the proof. ∎
See 1
Proof.
Let $P$ be a revealing POMDP with initial belief $b\in\Delta(\mathcal{S})$ and set of target states $\mathcal{X}\subseteq\mathcal{S}$. By Lemma 8, there exists a $(n,q)$-stopping optimal policy $\sigma$ with $n=|\mathcal{S}|+2$ and $q=\delta_{\min}^{2}\left(\tfrac{\delta_{\min}}{|\mathcal{S}|}\right)^{|\mathcal{S}|}$. By Lemma 10, for every $\varepsilon>0$, there exists a time horizon $T=\left\lceil\tfrac{n\log(\varepsilon)}{\log(1-q)}\right\rceil$ such that
$\mathbb{P}_{b}^{\sigma}(\exists t\leq T\quad B_{t}\in\mathcal{D}_{\mathcal{X}})\geq\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b)-\varepsilon.$
Moreover, by Lemma 11, we can compute an approximation of $\operatorname{val}_{\textnormal{BR}(\mathcal{X})}^{\,T}(b)$ denoted $\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{\,T}$ with total time complexity
$O\left(\dfrac{T^{|\mathcal{S}|}|\mathcal{S}|^{|\mathcal{S}|+1}|\mathcal{A}||\mathcal{Z}|}{\varepsilon^{|\mathcal{S}|-1}}\right).$
We consider the grid $G_{k}$ as previously defined in the proof of Lemma 11 with $k=\left\lceil\tfrac{(T+1)|\mathcal{S}|}{\varepsilon}\right\rceil$. For every $b\in\Delta(\mathcal{S})$, there exists $\Pi_{k}(b)\in G_{k}$ such that by taking $T=\left\lceil\tfrac{n\log(\varepsilon)}{\log(1-q)}\right\rceil$,
$\left|\operatorname{val}(b)_{\text{BR}(\mathcal{X})}-\widetilde{\operatorname{val}}_{\text{BR}(\mathcal{X})}^{\,T}\left(\Pi_{k}(b)\right)\right|$
$\qquad\leq\left|\operatorname{val}(b)_{\text{BR}(\mathcal{X})}-\mathbb{P}_{b}^{\sigma}(\exists t\leq T\quad B_{i}\in\mathcal{D}_{\mathcal{X}})\right|+\left|\mathbb{P}_{b}^{\sigma}(\exists t\leq T\quad B_{t}\in\mathcal{D}_{\mathcal{X}})-\widetilde{\operatorname{val}}_{\text{BR}(\mathcal{X})}^{\,T}\left(\Pi_{k}(b)\right)\right|$ (Triangle inequality)
$\qquad\leq\varepsilon+\left|\mathbb{P}_{b}^{\sigma}(\exists t\leq T\quad B_{t}\in\mathcal{D}_{\mathcal{X}})-\widetilde{\operatorname{val}}_{\text{BR}(\mathcal{X})}^{\,T}\left(\Pi_{k}(b)\right)\right|$ (Lemma 10)
$\qquad\leq 2\varepsilon.$ (Lemma 11)
Let $L\in\mathbb{N}$ be the bit-length bound on transition probabilities so that $\delta_{\min}\geq 2^{-L}$. Since $q\in(0,1]$ and using that $\log(1-q)\leq-q$, we have that
$T$ $\leq n\left(\dfrac{\log(\varepsilon)}{\log(1-q)}+1\right)$
$\leq n\left(\dfrac{1}{q}\log\left(\dfrac{1}{\varepsilon}\right)+1\right)$
$\leq n\left(\left(\dfrac{|\mathcal{A}|}{\delta_{\min}}\right)^{|\mathcal{S}|}\dfrac{1}{\delta_{\min}^{2}}\log\left(\dfrac{1}{\varepsilon}\right)+1\right)$
$\leq n\left(\left(|\mathcal{A}|2^{L}\right)^{|\mathcal{S}|}2^{2L}\log\left(\dfrac{1}{\varepsilon}\right)+1\right),$
and thus, we deduce the following upper bound for the horizon
$T=O\left((|\mathcal{S}|+2)2^{L(|\mathcal{S}|+2)}|\mathcal{A}|^{|\mathcal{S}|}\log\left(\dfrac{1}{\varepsilon}\right)\right)$
and thus, $T^{|\mathcal{S}|}=O\left((|\mathcal{S}|+2)^{|\mathcal{S}|}2^{L|\mathcal{S}|(|\mathcal{S}|+2)}|\mathcal{A}|^{{|\mathcal{S}|}^{2}}\log(1/\varepsilon)^{|\mathcal{S}|}\right)$. Inserting this bound into the above complexity yields
$\textnormal{Total time complexity}=O\left(\dfrac{(|\mathcal{S}|+2)^{|\mathcal{S}|}|\mathcal{S}|^{|\mathcal{S}|+1}|\mathcal{Z}|2^{L|\mathcal{S}|(|\mathcal{S}|+2)}|\mathcal{A}|^{{|\mathcal{S}|}^{2}+1}\log(1/\varepsilon)^{|\mathcal{S}|}}{\varepsilon^{|\mathcal{S}|-1}}\right).$
Therefore, the total time complexity is upper bounded by $2^{\textnormal{poly}(L,|\mathcal{S}|,\log(|\mathcal{A}|),\log(|\mathcal{Z}|),\log(1/\varepsilon))}$, which concludes the proof. ∎
Appendix B Proofs of Section 5
See 13
Proof.
Consider a POMDP $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, its underlying MDP $M$ with states $\mathcal{S}\times\mathcal{Z}$, an end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})$ of $M$, and a policy $\sigma$ on $P$ such that $\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})>0$. By contradiction, assume that there exists a state $q\in\mathcal{Q}$, an action $a\in\mathcal{E}(q)$, and a signal $z\in\mathcal{Z}$ such that, for all policies $\widetilde{\sigma}$ on $P$, we have
$\mathbb{P}_{b}^{\widetilde{\sigma}}(\mathcal{I}(\rho)\subseteq\mathcal{U})<1\,$, (1)
where $b$ is the belief after starting with belief $\mathds{1}[q]$, playing action $a$, and receiving signal $z$. We show that $\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})=0$, which is a contradiction.
By definition of end components, because $\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})>0$, for every action $a\in\mathcal{E}(q)$, the policy $\sigma$ plays action $a$ infinitely often when the belief of the controller has positive probability on $q$. By Equation 1, for all policies $\widetilde{\sigma}$ on $P$, the set $\mathcal{S}\setminus\mathcal{Q}$ is reached with positive probability, i.e.,
$\mathbb{P}_{b}^{\widetilde{\sigma}}(\exists t\geq 0\quad S_{t}\in\mathcal{S}\setminus\mathcal{Q})>0\,.$ (2)
We claim that the set $\mathcal{S}\setminus\mathcal{Q}$ is in fact reached within $2^{|\mathcal{Q}|}$ steps with positive probability. Formally, for all policies $\widetilde{\sigma}$,
$\mathbb{P}_{b}^{\widetilde{\sigma}}(\exists t\leq 2^{|\mathcal{Q}|}\quad S_{t}\in\mathcal{S}\setminus\mathcal{Q})>0\,.$ (3)
Indeed, we show that, starting from $b$, all policies reach a belief at which every action leads outside of $\mathcal{Q}$ with positive probability. Consider a belief $\widetilde{b}\subseteq\Delta(\mathcal{Q})$. Then, an action $a\in\mathcal{A}$ is safe for $\widetilde{b}$ if $\mathbb{P}_{\widetilde{b}}^{a}(\operatorname{supp}(B_{1})\subseteq\mathcal{Q})=1$. Note that, if an action is safe for a belief, then it is safe for all beliefs with the same support. Also, starting from $b$, if the controller can force the dynamic to visit only beliefs whose support has some safe action, then taking these actions one constructs a policy $\widetilde{\sigma}$ such that $\mathbb{P}_{b}^{\widetilde{\sigma}}(\forall t\geq 0\quad S_{t}\in\mathcal{Q})=1$. By Equation 2, this is not the case, i.e., for all policies the dynamic must visit a belief whose support has no safe action. Moreover, for each policy, such a belief is reached in the first $(2^{|\mathcal{Q}|}-1)$ steps because there are only so many supports, which proves our claim.
Equation 3 implies that the probability of reaching the set $\mathcal{S}\setminus\mathcal{Q}$ is lower bounded, i.e., for all policies $\widetilde{\sigma}$,
$\mathbb{P}_{b}^{\widetilde{\sigma}}\left(\exists t\leq 2^{|\mathcal{Q}|}\quad S_{t}\in\mathcal{S}\setminus\mathcal{Q}\right)\;\geq\;\left(\delta_{\min}\right)^{2^{|\mathcal{Q}|}}\,.$
With this, we prove that $\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})=0$, which is a contradiction.
Indeed, because $\mathds{1}[q]$ is visited infinitely often, action $a$ played infinitely often, and therefore the belief $b$ is visited infinitely often. Conditioned on the event $\{\inf(\rho)=\mathcal{U}\}$, visits to the state $q$ followed by playing the action $a$ occur infinitely often under policy $\sigma$. The above inequality implies that some states $s\in\mathcal{S}\setminus\mathcal{Q}$ are reached with lower-bounded positive probability after such visits. Therefore, we get
$\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=\mathcal{U})=0\,$,
which contradicts with the assumption, and therefore, yields the result. ∎
See 14
Proof.
By Lemma 13, for all states $q\in\mathcal{Q}$, actions $a\in\mathcal{E}(q)$, and signals $z\in\mathcal{Z}$, there exists a policy $\sigma_{q,a,z}$ such that
$\mathbb{P}_{b}^{\sigma_{q,a,z}}(\mathcal{I}(\rho)\subseteq\mathcal{U})\;=\;1\,$,
where $b$ is the belief after starting with belief $\mathds{1}[q]$, playing action $a$, and receiving signal $z$. We now construct a policy $\sigma_{q}$ such that
$\mathbb{P}_{\mathds{1}[q]}^{\sigma_{q}}(\mathcal{I}(\rho)=\mathcal{U})\;=\;1\,.$
[…]
See 15
Proof.
Consider a revealing POMDP $P$ with the initial belief $b_{0}\in\Delta(\mathcal{S})$. Denote the set of states for which, if the initial belief were a Dirac on that state, then the parity condition can be satisfied almost-surely by $\mathcal{X}\coloneqq\left\{s\in\mathcal{S}:\exists\sigma\quad\mathbb{P}_{\mathds{1}[s]}^{\sigma}(\operatorname{Parity})=1\right\}$. We show that, for all policies $\sigma\in\Sigma$,
$\mathbb{P}_{b_{0}}^{\sigma}\bigl(\rho\in\operatorname{Parity}\bigr)=\mathbb{P}_{b_{0}}^{\sigma}\bigl(\exists t\geq 0\quad B_{t}\in\mathcal{D}_{\mathcal{X}}\bigr)\,$,
which implies the equality by maximizing over the policy.
Fix a policy $\sigma$. Denote $\mathcal{W}$ the set of all end components of the underlying MDP on $\mathcal{S}\times\mathcal{Z}$ that can be reached by $\sigma$. Formally,
$\mathcal{W}\coloneqq\{\mathcal{U}:\mathbb{P}_{b_{0}}^{\sigma}(\mathcal{I}(\rho)=U)>0\}\,.$
Fix an end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})\in\mathcal{W}$. Note that, if a play $\rho$ eventually remains in $U$ forever, i.e., $\mathcal{I}(\rho)=\mathcal{U}$, then, by definition of end component, all the states $q\in\mathcal{Q}$ are visited infinitely often. Therefore, if $\mathcal{I}(\rho)=\mathcal{U}$, then
$\rho\in\operatorname{Parity}\iff\min\{\operatorname{pri}(q):q\in\mathcal{Q}\}\textnormal{ is even.}$
Consequently, the parity condition being satisfied depends on the end component reached. Formally, for all end components $\mathcal{U}\in\mathcal{W}$,
$\mathbb{P}_{b_{0}}^{\sigma}\bigl(\rho\in\operatorname{Parity}\,\mid\,\mathcal{I}(\rho)=\mathcal{U}\bigr)\in\{0,1\}\,.$
Denote by $\mathcal{W}_{\textnormal{P}}\subseteq\mathcal{W}$ the set of end components where the parity condition is satisfied with probability $1$ conditioned on eventually remaining in the end component. Formally,
$\mathcal{W}_{\textnormal{P}}\coloneqq\left\{\mathcal{U}\in\mathcal{W}:\mathbb{P}_{b_{0}}^{\sigma}\bigl(\rho\in\operatorname{Parity}\,\mid\,\mathcal{I}(\rho)=\mathcal{U}\bigr)=1\right\}\,.$
Recall that, by Lemma 12, some end component must be reached. Fix an arbitrary end component $\mathcal{U}=(\mathcal{Q},\mathcal{E})\in\mathcal{W}_{\textnormal{P}}$. By Lemma 14, for all states $q\in\mathcal{Q}$, there exists an almost-sure winning policy for the parity objective when the initial belief is $\mathds{1}[q]$. Therefore, $q$ is an almost-sure winning parity state. In conclusion, we get that
[…]
which yields the result. ∎
See 3
Proof.
By (Belly et al. 2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is in EXPTIME. By Theorem 1, approximating the belief-reachability value to this set of states is in EXPTIME. By Lemma 15, this value coincides with the parity value and therefore we computed an approximation of the parity value in EXPTIME. ∎
See 4
Proof.
Let $\mathcal{S}_{\textnormal{P}}$ be the set of almost-sure winning parity states. By Lemma 15, the parity value coincides with the belief-reachability value to the set $\mathcal{S}_{\textnormal{P}}$. By Corollary 9, there exists an optimal belief-reachability policy $\sigma_{0}$. For all state $s\in\mathcal{S}_{\textnormal{P}}$, let $\sigma_{s}$ be the almost-sure winning parity policy when the initial belief is $\mathds{1}[s]$. We now construct the optimal policy $\sigma$ as follows. The optimal policy follows the policy $\sigma_{0}$ until it reaches a Dirac belief $\mathds{1}[s]$ where $s\in\mathcal{S}_{\textnormal{P}}$. Then, it switches to following the policy $\sigma_{s}$. The probability of satisfying the parity condition under the policy $\sigma$ is the belief-reachability value to the set $\mathcal{S}_{\textnormal{P}}$, which coincides with the parity value. This completes the proof. ∎
See 5
Proof.
Consider revealing POMDPs with parity objectives. By Theorem 4, optimal policies exist. Therefore, if a revealing POMDP with parity objectives is limit-sure winning, then it is almost-sure winning. The reverse direction follows by definition. Lastly, by (Belly et al. 2025, Theorem 3), the existence of an almost-sure policy is in EXPTIME-complete. Therefore, limit-sure analysis is also EXPTIME-complete because the two problems coincide. ∎
Appendix C Outline of Algorithms
This section outlines three algorithms for quantitative analysis: for the $T$-step belief-reachability, for the belief-reachability, and for the parity value, respectively.
Given a belief $b\in\Delta(\mathcal{S})$, an action $a\in\mathcal{A}$, and a signal $z\in\mathcal{Z}$, recall that the belief update is defined by
$\tau(b,a,z)(s^{\prime})\coloneqq\frac{\sum_{s\in\mathcal{S}}b(s)\delta(s,a)(s^{\prime},z)}{\sum_{s,s^{\prime}\in\mathcal{S}}b(s)\delta(s,a)(s^{\prime},z)}\,.$
Outline of Algorithm 1
Given a revealing POMDP $P$, a target set $\mathcal{X}\subseteq\mathcal{S}$, a time horizon $T\in\mathbb{N}$, and an additive error $\varepsilon>0$, the algorithm returns a value that is guaranteed to be within $\varepsilon$ of the $T$-step belief-reachability value. Intuitively, Algorithm 1 replaces the continuous set of beliefs with a finite uniform grid to run value iteration only on those grid points.
The algorithm has two stages and follows from the proof of Lemma 11:
  • $(1)$
    We build a finite uniform grid of beliefs $G_{k}\subseteq\Delta(\mathcal{S})$ such that every belief is at most $\varepsilon/(T+1)$ away from some grid point.
  • $(2)$
    We perform $T$ backward Bellman updates on the grid: each grid belief evaluates every action and keeps the one with the best expected value over the projected successor beliefs.
Because the $T$-step belief-reachability value is Lipschitz, the projection errors accumulated over the steps is at most $\varepsilon$. Therefore, the algorithm returns an epsilon approximation of the $T$-step belief-reachability value.
Outline of Algorithm 2
Given a revealing POMDP $P$, a target set $\mathcal{X}\subseteq\mathcal{S}$, and an additive error $\varepsilon>0$, the algorithm return a value that is guaranteed to be within $\varepsilon$ of the belief-reachability value. Intuitively, Algorithm 2 approximates the belief-reachability value by computing the $T$-step belief reachability value for a horizon $T$ chosen large enough that the two differ by at most $\varepsilon$.
The algorithm proceeds as follows:
  • $(1)$
    We first choose $T$ large enough. By Lemma 7 and Lemma 8, there is an $(n,q)$-stopping optimal policy $\sigma$ with parameters $n=|\mathcal{S}|+2$ and $q=\delta_{\min}^{2}(\delta_{\min}/|\mathcal{A}|)^{|\mathcal{S}|}$ for the belief-reachability objectives. Next, by Lemma 10, we set $T=n\lceil\log(\varepsilon/2)/\log(1-q)\rceil$ so that the $T$-step belief-reachability objective under $\sigma$ is an $\varepsilon/2$ approximation of the belief-reachability value.
  • $(2)$
    We call Lemma 11 by taking $T=n\lceil\log(\varepsilon/2)/\log(1-q)\rceil$ and $\varepsilon/2$.
By Theorem 1, the algorithm returns an epsilon approximation of the belief-reachability value.
Outline of Algorithm 3
Given a revealing POMDP $P$, a priority function $\operatorname{pri}:\mathcal{S}\to\{0,\dots,d\}$, and an additive error $\varepsilon>0$, the algorithm returns a value that is guaranteed to be within $\varepsilon$ of the parity value. Intuitively, Algorithm 1 computes the approximation of the belief-reachability objectives to the set of Dirac on the almost-sure winning parity states.
The algorithm proceeds as follows:
  • $(1)$
    We use (Belly et al. 2025, Theorem 3) to compute the set of almost-sure winning states, denoted $\mathcal{S}_{P}$.
  • $(2)$
    We call Algorithm 2 by taking the target set as $\mathcal{S}_{P}$.
By Lemma 15, the parity value of $P$ and the belief-reachability value to the set of Dirac on the almost-sure winning parity states objective coincide. Therefore, because Algorithm 2 returns an epsilon approximation of the belief reachability value, the algorithm returns an epsilon approximation of the parity value.
Input: Revealing POMDP $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, target set $\mathcal{X}\subseteq\mathcal{S}$, horizon $T\in\mathbb{N}$, and additive error $\varepsilon>0$
Output: $v$ such that $\left|v-\operatorname{val}^{T}_{\textnormal{BR}(\mathcal{X})}(b_{0})\right|\leq\varepsilon$
1
[…]
3
[…]
5
6foreach $b\in G_{k}$ do
7 $\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{0}(b)\leftarrow\begin{cases}1&\text{if }b\in\mathcal{D}_{\mathcal{X}},\\ 0&\text{otherwise.}\end{cases}$;
8
9 end foreach
10
11for $t\leftarrow 1$ to $T$ do
12 foreach $b\in G_{k}$ do
13    if $b\in\mathcal{D}_{\mathcal{X}}$ then $\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t}(b)\leftarrow 1$;
14    else foreach $a\in\mathcal{A}$ do
15       $Q_{a}\leftarrow 0$ ;
16       foreach $z\in\mathcal{Z}$ do
17          foreach $s,s^{\prime}\in\mathcal{S}$ do
18             $p\leftarrow b(s)\,\delta(s,a)(s^{\prime},z)$;
19             if $p>0$ then
20                $b^{\prime}\leftarrow\Pi_{k}(\tau(b,a,z))$;
21                $Q_{a}\leftarrow Q_{a}+p\cdot\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t-1}(b^{\prime})$;
22
23               end if
24
25            end foreach
26
27         end foreach
28
29      end foreach
30    $\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{t}(b)\leftarrow\max_{a\in A}Q_{a}$ ;
31
32   end foreach
33
34 end for
35return $\widetilde{\operatorname{val}}_{\textnormal{BR}(\mathcal{X})}^{T}(\Pi_{k}(b))$;
Algorithm 1 Approximation of $T$-step belief-reachability value
Input: Revealing POMDP $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, target set $\mathcal{X}\subseteq\mathcal{S}$, and additive error $\varepsilon>0$
Output: $v$ such that $\left|v-\operatorname{val}_{\textnormal{BR}(\mathcal{X})}(b_{0})\right|\leq\varepsilon$
1
[…]
3
[…]
5
[…]
7
[…]
9
10return $v$;
Algorithm 2 Approximation of belief-reachability value
Input: Revealing POMDP $P=(\mathcal{S},\mathcal{A},\mathcal{Z},\delta,b_{0})$, priority function $\operatorname{pri}:\mathcal{S}\to\{0,\dots,d\}$, and additive error $\varepsilon>0$
Output: $v$ such that $\left|v-\operatorname{val}_{\textnormal{P}}(b_{0})\right|\leq\varepsilon$
1
[…]
3
[…]
5
6return $v$;
Algorithm 3 Approximation of parity value

### v2

@@
 POMDPs
-Partially observable Markov decision processes (POMDPs) model sequential decision-making under uncertainty (Bertsekas 1976; Papadimitriou and Tsitsiklis 1987; Kaelbling, Littman, and Cassandra 1998).
+Partially observable Markov decision processes (POMDPs) model sequential decision-making under uncertainty Bertsekas (1976); Papadimitriou and Tsitsiklis (1987); Kaelbling, Littman, and Cassandra (1998).
 At each step, the environment is in some hidden state.
@@
 The controller cannot observe the state directly, but observes a signal that discloses only partial information of the state.
-POMDPs generalize classic models: Markov decision processes (MDPs), in which the state is fully observed (Puterman 2014), and blind MDPs, in which no state information is observed and are equivalent to Probabilistic Finite Automata (Rabin 1963; Paz 1971).
+POMDPs generalize classic models: Markov decision processes (MDPs), in which the state is fully observed Puterman (2014), and blind MDPs, in which no state information is observed and are equivalent to Probabilistic Finite Automata Rabin (1963); Paz (1971).
 Objectives
 The controller aims to maximize an objective function, which formally captures the desired behaviors of the model.
-Two main classes of objectives are typically considered: logical objectives, e.g., reachability and LTL (linear-temporal logic), and quantitative objectives, e.g., discounted sum and limit-average; see (Puterman 2014; Baier and Katoen 2008; Filar and Vrieze 1997) for details.
+Two main classes of objectives are typically considered: logical objectives, e.g., reachability and LTL (linear-temporal logic), and quantitative objectives, e.g., discounted sum and limit-average; see Puterman (2014); Baier and Katoen (2008); Filar and Vrieze (1997) for details.
 This work focuses on logical objectives.
@@
 The objective is satisfied if the smallest priority appearing infinitely often is even.
-Parity objectives express all commonly used temporal objectives such as liveness, and are a canonical form to express all $\omega$-regular objectives (Thomas 1997) and LTL objectives (Baier and Katoen 2008).
+Parity objectives express all commonly used temporal objectives such as liveness, and are a canonical form to express all $\omega$-regular objectives Thomas (1997) and LTL objectives Baier and Katoen (2008).
 Hence, the study of POMDPs with parity objectives is a fundamental theoretical problem.
 Applications
-POMDPs have been applied across diverse fields, including computational biology (Durbin et al.
-1998) and reinforcement learning (Kaelbling, Littman, and Moore 1996).
-In particular, POMDPs with logical objectives have proven useful in many application areas such as probabilistic planning (Bonet and Geffner 2009); randomized embedded scheduler (de Alfaro et al.
-2005); randomized distributed algorithms (Pogosyants, Segala, and Lynch 2000); probabilistic specification languages (Baier, Größer, and Bertrand 2012); and robot planning (Kress-Gazit, Fainekos, and Pappas 2009; Chatterjee et al.
-2015).
+POMDPs have been applied across diverse fields, including computational biology Durbin et al.
+(1998) and reinforcement learning Kaelbling, Littman, and Moore (1996).
+In particular, POMDPs with logical objectives have proven useful in many application areas such as probabilistic planning Bonet and Geffner (2009); randomized embedded scheduler de Alfaro et al.
+(2005); randomized distributed algorithms Pogosyants, Segala, and Lynch (2000); probabilistic specification languages Baier, Größer, and Bertrand (2012); and robot planning Kress-Gazit, Fainekos, and Pappas (2009); Chatterjee et al.
+(2015).
 Computational Analysis Questions
@@
 Most results for POMDPs and the computational analysis are negative (undecidability results).
-Indeed, the limit-sure analysis for reachability objectives is undecidable (Gimbert and Oualhadj 2010; Chatterjee and Henzinger 2010), which extends to general parity objectives.
-The almost-sure analysis for reachability objectives is EXPTIME-complete (Baier, Bertrand, and Größer 2008; Chatterjee, Doyen, and Henzinger 2010), but is undecidable for parity objectives even for two priorities (namely coBüchi objectives) (Baier, Bertrand, and Größer 2008; Chatterjee et al.
-2010).
-The quantitative analysis for reachability objectives is undecidable (Madani, Hanks, and Condon 2003), which extends to general parity objectives.
+Indeed, the limit-sure analysis for reachability objectives is undecidable Gimbert and Oualhadj (2010); Chatterjee and Henzinger (2010), which extends to general parity objectives.
+The almost-sure analysis for reachability objectives is EXPTIME-complete Baier, Bertrand, and Größer (2008); Chatterjee, Doyen, and Henzinger (2010), but is undecidable for parity objectives even for two priorities (namely coBüchi objectives) Baier, Bertrand, and Größer (2008); Chatterjee et al.
+(2010).
+The quantitative analysis for reachability objectives is undecidable Madani, Hanks, and Condon (2003), which extends to general parity objectives.
 Given this wide range of results about non-existence of algorithms a natural direction to explore is the existence of subclasses for which the computational problems are decidable.
@@
 the probability distribution over states (the belief), occasionally collapses into a Dirac distribution on the announced state.
-Revealing POMDPs have been previously studied in several works, including (Chen and Liew 2023; Belly et al.
-2025; Avrachenkov, Dhiman, and Kavitha 2025), which present motivation and applications for this model.
+Revealing POMDPs have been previously studied in several works, including Chen and Liew (2023); Belly et al.
+(2025); Avrachenkov, Dhiman, and Kavitha (2025), which present motivation and applications for this model.
 Previous Results and Open Questions
-A key result for revealing POMDPs shows that the almost-sure analysis for parity objectives is EXPTIME-complete (Belly et al.
-2025).
+A key result for revealing POMDPs shows that the almost-sure analysis for parity objectives is EXPTIME-complete Belly et al.
+(2025).
 However, the limit-sure and quantitative analysis problems for reachability and parity objectives remained open for revealing POMDPs.
@@
 Technical Contributions
-A closely related work established that almost-sure analysis for parity objectives in revealing POMDPs is EXPTIME-complete (Belly et al.
-2025).
-Additionally, (Chatterjee, Chmelík, and Tracol 2016) previously showed that almost-sure analysis for parity objectives in general POMDPs under finite-memory policies is EXPTIME-complete.
+A closely related work established that almost-sure analysis for parity objectives in revealing POMDPs is EXPTIME-complete Belly et al.
+(2025).
+Additionally, Chatterjee, Chmelík, and Tracol (2016) previously showed that almost-sure analysis for parity objectives in general POMDPs under finite-memory policies is EXPTIME-complete.
 Therefore, the EXPTIME-completeness result naturally follows once finite memory policies are proven sufficient for almost-sure analysis.
-However, limit-sure analysis and quantitative analysis for POMDPs remain undecidable in general, even under finite memory policies (Chatterjee, Saona, and Ziliotto 2021).
+However, limit-sure analysis and quantitative analysis for POMDPs remain undecidable in general, even under finite memory policies Chatterjee, Saona, and Ziliotto (2021).
 This highlights a sharp contrast with the almost-sure analysis and gives rise to new technical challenges.
@@
 The study of subclasses of POMDPs with tractable algorithms is a broad topic with many directions.
-For example, subclasses of POMDPs for qualitative analysis have been explored in (Fijalkow et al.
-2015; Chatterjee and Tracol 2012); and for quantitative analysis various subclasses have also been studied such as with ergodicity condition (Chatterjee et al.
-2024); multiple environments only (Van Der Vegt, Jansen, and Junges 2023; Chatterjee et al.
-2025); or in online learning (Liu et al.
-2022; Chen et al.
-2023).
+For example, subclasses of POMDPs for qualitative analysis have been explored in Fijalkow et al.
+(2015); Chatterjee and Tracol (2012); and for quantitative analysis various subclasses have also been studied such as with ergodicity condition Chatterjee et al.
+(2024); multiple environments only Van Der Vegt, Jansen, and Junges (2023); Chatterjee et al.
+(2025); or in online learning Liu et al.
+(2022); Chen et al.
+(2023).
 Our work focuses on such a subclass, namely revealing POMDPs, which has been studied in the literature.
@@
 They considered an application of strongly-connected revealing blind MDPs.
-(Chen and Liew 2023) introduced the model of revealing blind MDPs.
+Chen and Liew (2023) introduced the model of revealing blind MDPs.
 They consider the discounted objective and present algorithms based on finite-horizon approximation to compute the discounted value.
@@
   • $b_{0}\in\Delta(\mathcal{S})$ is the initial belief over the states.
-Markov Decision Processes (MDPs) are POMDPs in which the signal corresponds to the state (Puterman 2014).
+Markov Decision Processes (MDPs) are POMDPs in which the signal corresponds to the state Puterman (2014).
 Formally, an MDP is denoted by $M=(\mathcal{S},\mathcal{A},\delta,b_{0})$ with transition function $\delta\colon\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$.
@@
 $\qquad\implies\sum_{\widetilde{s}\in\mathcal{S}}\delta(s,a)(\widetilde{s},s^{\prime})=\delta(s,a)(s^{\prime},s^{\prime})>0\,.$
-Revealing POMDPs coincide with the class of strongly revealing defined in (Belly et al.
-2025).
+Revealing POMDPs coincide with the class of strongly revealing defined in Belly et al.
+(2025).
 Indeed, if the controller observes a signal $s\in\mathcal{Z}$, then their next belief is $\mathds{1}[s]$, which corresponds to a revelation of the state.
@@
 Even with exponentially many signals encoded in a polynomial size input, our EXPTIME upper bound holds.
-  • In general POMDPs, it is enough to consider transition functions that pair each state with only one signal, known as deterministic observation, see (Chatterjee, Chmelík, and Tracol 2016, Remark 4).
+  • In general POMDPs, it is enough to consider transition functions that pair each state with only one signal, known as deterministic observation, see (Chatterjee, Chmelík, and Tracol, 2016, Remark 4).
 For revealing POMDPs, deterministic observations do not capture the general case because they correspond to MDPs.
@@
 For a finite prefix of a play, an element in $(\mathcal{S}\times\mathcal{A})^{*}$, its cone is the set of plays with it as their prefix.
-Given a policy $\sigma$ and an initial belief $b_{0}$, the unique probability measure over Borel sets of infinite plays obtained given $\sigma$ is denoted by $\mathbb{P}^{\sigma}_{b_{0}}(\cdot)$, which is defined by Carathéodory’s extension theorem by extending the natural definition over cones of plays (Billingsley 2012).
+Given a policy $\sigma$ and an initial belief $b_{0}$, the unique probability measure over Borel sets of infinite plays obtained given $\sigma$ is denoted by $\mathbb{P}^{\sigma}_{b_{0}}(\cdot)$, which is defined by Carathéodory’s extension theorem by extending the natural definition over cones of plays Billingsley (2012).
 Objectives
-An objective in a POMDP is a Borel set of plays $\Phi\subseteq\Omega$ in the Cantor topology on $\Omega$ (Kechris 1995).
-We consider objectives in the first $2\nicefrac{{1}}{{2}}$ levels of the Borel hierarchy, including the parity objective which expresses all $\omega$-regular objectives (Thomas 1997).
+An objective in a POMDP is a Borel set of plays $\Phi\subseteq\Omega$ in the Cantor topology on $\Omega$ Kechris (1995).
+We consider objectives in the first $2\nicefrac{{1}}{{2}}$ levels of the Borel hierarchy, including the parity objective which expresses all $\omega$-regular objectives Thomas (1997).
 Denote the state at time $t$ by $s_{t}$ and set of states that occur infinitely often in a play $\rho$ by $\mathcal{I}(\rho)\coloneqq\{s\in\mathcal{S}:\forall t\geq 0\;\exists\widetilde{t}\geq t\quad s=s_{\widetilde{t}}\in\rho\}$.
@@
   • Reachability: Given a set $\mathcal{X}\subseteq\mathcal{S}$ of target states, the reachability objective requires that a target state is visited at least once.
-Formally, the reachability objective is $\operatorname{Reach}(\mathcal{X})\coloneqq\{\rho\in\Omega:\exists t\geq 0\quad s_{t}\in\mathcal{X}\}$.
+Formally, the reachability objective is
+    $\operatorname{Reach}(\mathcal{X})\coloneqq\{\rho\in\Omega:\exists t\geq 0\quad s_{t}\in\mathcal{X}\}.$
   • Parity: Given a $d\geq 0$ and a function $\operatorname{pri}\colon\mathcal{S}\to\{0,1,\ldots,d\}$ of priorities, the parity objective requires that the smallest priority that appears infinitely often is even.
-Formally, $\operatorname{Parity}\coloneqq\left\{\rho\in\Omega:\min\{\operatorname{pri}(s):s\in\mathcal{I}(\rho)\}\text{ is even}\right\}$.
+Formally,
+    $\operatorname{Parity}\coloneqq\left\{\rho\in\Omega:\min\{\operatorname{pri}(s):s\in\mathcal{I}(\rho)\}\text{ is even}\right\}.$
 Value
@@
 Optimal policies exist for parity objectives for revealing POMDPs.
-The following results follow from (Belly et al. 2025, Theorem 3) and Theorem 4.
+The following results follow from (Belly et al., 2025, Theorem 3) and Theorem 4.
 Corollary 5.
@@
 We say that a policy is stopping if it is stopping for some $n$ and $q>0$.
-Similar to (Andersson and Miltersen 2009, Lemma 5) that considers finite duration games, it is easy to see that, if a policy is stopping and uses only reliable actions, then it is optimal.
+Similar to (Andersson and Miltersen, 2009, Lemma 5) that considers finite duration games, it is easy to see that, if a policy is stopping and uses only reliable actions, then it is optimal.
 Lemma 7.
@@
 Denote $\delta_{\min}$ the minimum non-zero probability in the transition function, i.e.,
-$\delta_{\min}\coloneqq\min\Bigl\{\delta(s,a)(s^{\prime},z)\;:\;\forall s,\,s^{\prime}\in\mathcal{S},\;a\in\mathcal{A},\;z\in\mathcal{Z}$
-$\delta(s,a)(s^{\prime},z)>0\Bigr\}.$
+$\delta_{\min}\coloneqq\min\Bigl\{\delta(s,a)(s^{\prime},z)\;:\;\forall s,\,s^{\prime}\in\mathcal{S},\;a\in\mathcal{A},\;z\in\mathcal{Z}\quad\delta(s,a)(s^{\prime},z)>0\Bigr\}.$
 Lemma 8.
@@
 This approach implies a 2EXPTIME complexity upper bound.
-A fundamental result states that, for horizons that are polynomially large with respect to the input, computing the finite horizon value of POMDPs is PSPACE-complete (Papadimitriou and Tsitsiklis 1987, Theorem 6, page 448).
+A fundamental result states that, for horizons that are polynomially large with respect to the input, computing the finite horizon value of POMDPs is PSPACE-complete (Papadimitriou and Tsitsiklis, 1987, Theorem 6, page 448).
 The technique, instead of listing explicitly all exponentially many histories, compactly represents them using nondeterminism and space proportional to the horizon.
@@
 Instead, we prove an EXPTIME upper bound.
-Point-based algorithms were introduced for POMDPs by (Pineau, Gordon, and Thrun 2003) as an alternative to approximating the value of POMDPs by considering a fixed subset of beliefs and projected belief updates.
-For a survey on point-based algorithms see (Shani, Pineau, and Kaplow 2013; Walraven and Spaan 2019), which focuses on finite horizon objectives but does not provide the complexity upper bound we require.
+Point-based algorithms were introduced for POMDPs by Pineau, Gordon, and Thrun (2003) as an alternative to approximating the value of POMDPs by considering a fixed subset of beliefs and projected belief updates.
+For a survey on point-based algorithms see Shani, Pineau, and Kaplow (2013); Walraven and Spaan (2019), which focuses on finite horizon objectives but does not provide the complexity upper bound we require.
 Finite-Horizon Belief-Reachability
@@
 In the sequel, we show that, for revealing POMDPs, the parity objective value coincides with the belief-reachability to the set of Dirac on the almost‐sure winning parity states.
-This proof uses the standard notion of end components in MDPs, introduced in (Alfaro 1998); see also (Baier and Katoen 2008, Section 10.6.3).
+This proof uses the standard notion of end components in MDPs, introduced in Alfaro (1998); see also (Baier and Katoen, 2008, Section 10.6.3).
 Underlying MDPs of POMDPs
@@
 The following statement is a fundamental result of end components in MDPs.
-Lemma 12 ((Alfaro 1998, Theorems 3.1 and 3.2)).
+Lemma 12 ((Alfaro, 1998, Theorems 3.1 and 3.2)).
 For MDPs, the following assertions hold.
@@
 Example 1.
-Consider the example presented in (Chatterjee, Saona, and Ziliotto 2021, Example 4.3) of a POMDP with only one possible signal in Figure 1, commonly known as a blind MDP.
+Consider the example presented in (Chatterjee, Saona, and Ziliotto, 2021, Example 4.3) of a POMDP with only one possible signal in Figure 1, commonly known as a blind MDP.
 It has four states $\mathcal{S}=\{s_{0},s_{1},\bot,\top\}$ and two actions: wait ($w$) and commit ($c$), i.e., $\mathcal{A}=\{w,c\}$.
@@
 Proof Sketch.
-By (Belly et al. 2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is EXPTIME. Therefore, this result is a direct implication of Lemma 15 and Theorem 1. ∎
+By (Belly et al., 2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is EXPTIME. Therefore, this result is a direct implication of Lemma 15 and Theorem 1. ∎
 See 4
@@
 Proof.
-By (Belly et al.
-2025, Theorem 3), computing almost-sure winning parity states in the revealing POMDP is in EXPTIME.
+By [7, Theorem 3], computing almost-sure winning parity states in the revealing POMDP is in EXPTIME.
 By Theorem 1, approximating the belief-reachability value to this set of states is in EXPTIME.
@@
 The reverse direction follows by definition.
-Lastly, by (Belly et al.
-2025, Theorem 3), the existence of an almost-sure policy is in EXPTIME-complete.
+Lastly, by [7, Theorem 3], the existence of an almost-sure policy is in EXPTIME-complete.
 Therefore, limit-sure analysis is also EXPTIME-complete because the two problems coincide.
@@
   • $(1)$
-    We use (Belly et al. 2025, Theorem 3) to compute the set of almost-sure winning states, denoted $\mathcal{S}_{P}$.
+    We use [7, Theorem 3] to compute the set of almost-sure winning states, denoted $\mathcal{S}_{P}$.
   • $(2)$